### v1

# PriceFM: Foundation Model for Probabilistic Electricity Price Forecasting

Abstract
Electricity price forecasting in Europe presents unique challenges due to the continent’s increasingly integrated and physically interconnected power market. While recent advances in deep learning and foundation models have led to substantial improvements in general time series forecasting, most existing approaches fail to capture the complex spatial interdependencies and uncertainty inherent in electricity markets. In this paper, we address these limitations by introducing a comprehensive and up-to-date dataset across 24 European countries (38 regions), spanning from 2022-01-01 to 2025-01-01. Building on this groundwork, we propose PriceFM, a spatiotemporal foundation model that integrates graph-based inductive biases to capture spatial interdependencies across interconnected electricity markets. The model is designed for multi-region, multi-timestep, and multi-quantile probabilistic electricity price forecasting. Extensive experiments and ablation studies confirm the model’s effectiveness, consistently outperforming competitive baselines and highlighting the importance of spatial context in electricity markets. The dataset and code can be found at https://github.com/runyao-yu/PriceFM.
Introduction
The European electricity market is physically interconnected through a network of cross-border transmission lines, enabling the exchange of electricity between regions and optimizing the social welfare at the European level (Lago et al. 2018). However, physical constraints, such as limited transmission capacity, can restrict electricity flow between regions and lead to zonal price differences (Finck 2021), illustrated in Figure 1. These price disparities highlight the spatial nature of electricity price formation. Recent studies show that electricity price dynamics are strongly influenced by spatial interdependencies and cannot be accurately captured using region-specific models (Do et al. 2024). Therefore, explicitly modeling the spatial structure of the European electricity market is essential for producing accurate price forecasts.
Figure 1: Spatial distribution of electricity price and number of neighboring regions. a Electricity prices for 38 European regions averaged from 2022-01-01 to 2025-01-01. A significant zonal price difference is observed between north and south regions. b Number of neighboring regions that are directly connected to certain region via transmission lines. For example, France (FR) and Portugal (PT) are directly connected to Spain (ES), thus the number of regions for ES is 2.
Most existing studies on electricity price forecasting do not explicitly model the spatial structure and focus on a single-region market, particularly Germany (Muniain and Ziel 2020; Maciejowska, Nitka, and Weron 2021; Kitsatoglou et al. 2024), as the German market is one of the largest markets in Europe. Other studies explore forecasting methods for markets such as Denmark, Finland, and Spain, also using region-specific models (Ziel and Weron 2018; Gianfreda, Ravazzolo, and Rossini 2020; Loizidis, Kyprianou, and Georghiou 2024). More recent works explicitly model the spatial nature of the electricity price. For instance, a Graph Convolutional Network (GCN) is applied to capture spatial interdependencies in the Nordic markets, such as Norway, Sweden, and Finland (Yang et al. 2024). Moreover, an attention-based variant is developed to predict prices in certain European markets such as Austria, Germany, and Hungary (Meng et al. 2024). However, these models cover only subsets of Europe and primarily produce point forecasts, failing to capture the uncertainty inherent in electricity prices.
Uncertainty modeling in electricity markets is critical, as the electricity price is strongly influenced by intermittent renewable generation and fluctuating demand (Lago et al. 2021). Consequently, price forecasting should extend beyond traditional pointwise forecasting to explicitly quantify uncertainties, especially for applications involving risk-sensitive applications such as energy trading and operational planning (Ziel and Steinert 2018). A comprehensive survey summarizes various probabilistic forecasting approaches based on quantile regression methods (Lago et al. 2021). However, these existing methods often focus on single-region markets, thereby neglecting the rich spatial information.
In recent years, foundation models for time series forecasting have achieved remarkable success across diverse domains, demonstrating strong generalization capabilities by capturing complex data patterns (Zhou et al. 2022; Wu et al. 2023; Wang et al. 2024), which makes them attractive candidates for electricity price forecasting. However, existing models are primarily designed for general univariate or multivariate time series tasks and are not explicitly tailored to electricity markets (Liang et al. 2024). Moreover, some foundation models only provide pointwise forecasts and thus fail to capture the uncertainty essential for risk-aware trading decisions in volatile markets. Therefore, addressing the spatial interdependencies and uncertainty in electricity price forecasting within the integrated European market necessitates a tailored foundation model.
To support the development of foundation models for electricity price forecasting, there is a pressing need for high-quality, large-scale, and up-to-date datasets that reflect the spatiotemporal complexity of the integrated European market. However, existing datasets are often fragmented in structure, cover only short time periods, are outdated, or focus on individual regions (Lago et al. 2021). This lack of standardized, cross-border data poses a significant barrier to training and evaluating foundation models.
In this paper, we bridge the gap between recent advances in foundation models and the requirements of electricity price forecasting in spatially interconnected markets. To this end, we introduce a comprehensive and up-to-date dataset capturing key spatiotemporal drivers of electricity prices across Europe. As a step forward, we propose PriceFM, a foundation model that incorporates graph-based inductive biases to learn spatial structures and generate probabilistic forecasts tailored to the integrated European electricity market.
Contribution
  • We introduce and release a comprehensive, up-to-date dataset. To the best of our knowledge, this is the largest and most diverse open dataset for European electricity markets, comprising day-ahead electricity prices, load forecasts, and solar and wind power generation forecasts (onshore and offshore), covering 24 European countries (38 regions), spanning from 2022-01-01 to 2025-01-01.
  • We propose and release the PriceFM, a novel forecasting framework that integrates prior graph knowledge derived from the spatial topology of the European electricity market. PriceFM supports multi-region, multi-timestep, and multi-quantile probabilistic electricity price forecasting.
  • We conduct extensive experiments to evaluate the model’s performance against multiple baselines, and assess the impact of design choices through ablation studies, thereby providing both quantitative evidence of overall performance and empirical insights into optimal configurations.
Preliminary
The forecasting target is $\mathcal{T}=24$ hourly prices for the delivery day $\mathcal{D}+1$, using data available before gate closure, typically around midday on day $\mathcal{D}$. After midday on $\mathcal{D}$, the electricity prices for $\mathcal{D}+1$ are published and known. We employ a backward-looking window of size L{L} (e.g. L=24{L}=24 corresponds to 24 hours from $\mathcal{D}$), for known electricity prices, denoted as $\mathbf{X}^{\mathrm{price}}_{r_{\mathrm{in}}}$. We also include forward-looking exogenous features, such as day‑ahead forecasts of load, solar, and wind (onshore and offshore) power generation for $\mathcal{D}+1$, denoted as $\mathbf{X}^{\mathrm{exo}}_{r_{\mathrm{in}}}$, made on $\mathcal{D}$ before gate closure, as well as their historical values over L{L}. The forecasting setup and the choice of feature set are widely used in prior works (Maciejowska 2020; Uniejewski and Weron 2021; Meng et al. 2024), illustrated in Figure 2. Note that this work aims to utilize multi-region inputs to produce multi-region, multi-timestep, and multi-quantile probabilistic electricity price forecasts. Therefore, the input and output are defined as:
  • Input: $\mathbf{X}^{\mathrm{price}}_{r_{\mathrm{in}}}\;\in\;\mathbb{R}^{L\times f_{1}}$ and $\mathbf{X}^{\mathrm{exo}}_{r_{\mathrm{in}}}\;\in\;\mathbb{R}^{(L+\mathcal{T})\times f_{2}}.$
  • Output: $\hat{\mathbf{y}}_{{r_{\mathrm{out}}},\tau}\in\mathbb{R}^{\mathcal{T}}$
where $r_{\mathrm{in}},r_{\mathrm{out}}\in\mathcal{R}=\{\text{AT},\dots,\text{SK}\}$ (region codes, detailed in Appendix, Table 1.), $\tau\in\mathcal{Q}=\{0.1,0.5,0.9\}$ (quantile levels), $f_{1}=1$, and $f_{2}$ varies by region. For example, the forecasted wind power generation is not available in SI.
Figure 2: Forecasting setup and feature set.
Figure 3: European-level energy data in 2024, averaged across all available regions (depending on data availability). a Electricity price. Price spikes sharply during the morning and evening peak, dip around midday, and shows higher volatility in winter. b Forecasted load. Load exhibits a clear double‐peak each day (in morning and evening) with winter peaks substantially larger than summer. c Forecasted solar power generation. Solar is zero overnight, rises in a smooth bell curve to a strong midday maximum, then falls back to zero by dusk, and is much higher in summer. d Forecasted wind power generation (sum of onshore and offshore). Wind lacks a daily pattern, fluctuates with high‐frequency spikes, and is much higher in the windy winter.
Data
The dataset is primarily collected from the European Network of Transmission System Operators for Electricity (ENTSO-E). We release the raw and cleaned dataset.
Spatiotemporal Coverage
Spatially, the dataset covers 24 European countries (38 regions). These regions reflect transmission zones rather than administrative boundaries. For example, DK is split into two regions: DK1 and DK2. Each is connected to different neighboring regions, resulting in distinct cross-border power flows. Temporally, the dataset spans from 2022-01-01 to 2025-01-01, providing wide temporal coverage.
Feature Set
The feature set includes day-ahead electricity prices, load forecasts, and solar and wind power generation forecasts (onshore and offshore). For simplicity, we refer to these features as price, load, solar, and wind (onshore and offshore), respectively. Importantly, some regions only contain partial features. The full availability of features across regions is detailed in Appendix, Table 1. A European-level visualization of these features is shown in Figure 3.
Resolution
We resample all features in an hourly resolution, as the raw data exhibit a heterogeneous temporal structure. For example, load from ES is provided at an hourly resolution before 2022-05-23 and then switches to a quarter-hourly resolution afterward; the price from AT is reported hourly, while the load is reported quarter-hourly.
Missing Value
Partial features are excluded due to the high rate (above 15%) of missing values, summarized in Appendix, Table 1. For example, wind offshore from FR has a missing rate of 53.2% and is only available after 2023-08-07; solar from LV has a 75.5% missing rate and is only available after 2024-04-07; load from SK has a missing rate of 16.8% and is no longer available after 2024-07-01. For features with low missing rates (around 1%), such as load from EE and SI, we fill them using linear interpolation.
Figure 4: Structure of PriceFM. The input features $\mathbf{X}^{\mathrm{price}}_{r_{\mathrm{in}}}$ and $\mathbf{X}^{\mathrm{exo}}_{r_{\mathrm{in}}}$ are passed into a fusion block to learn regional representation $\mathbf{F}_{r_{\mathrm{in}}}$. These regional representations are then stacked to form the spatial representation $\mathbf{S}$. Next, $\mathbf{S}$ is passed to the graph block to produce the representation $\mathbf{U}_{r_{\mathrm{out}}}$, which incorporates prior graph knowledge. Finally, these prior-knowledge-injected representations are fed into hierarchical quantile heads to produce multi-region, multi-timestep, and multi-quantile predictions.
Model
PriceFM, illustrated in Figure 4, aims to produce joint forecasts for multi-region, multi-timestep, and multi-quantile settings. The inputs $\mathbf{X}^{\mathrm{price}}_{r_{\mathrm{in}}}\in\mathbb{R}^{L\times f_{1}}$ and $\mathbf{X}^{\mathrm{exo}}_{r_{\mathrm{in}}}\in\mathbb{R}^{(L+\mathcal{T})\times f_{2}}$ for input region $r_{\mathrm{in}}$} are described in Section Preliminary. The inclusion of price and exogenous features introduces heterogeneity along the temporal and feature dimensions.
Backbone
We design a Fusion Block to enable effective fusion of heterogeneous multivariate input features and produce a spatial representation, and a Graph Block to incorporate prior graph knowledge to enforce spatial decay.
Fusion Block
We first align the temporal dimensions by projecting the price feature to length $L+\mathcal{T}$ via a dense layer with linear activation, followed by a transformation into a hidden space of dimension $h$ via $k$ dense layers:
$\mathbf{X}^{\mathrm{price}}_{r_{\mathrm{in}}}$ $\xrightarrow{\mathrm{Project}}\tilde{\mathbf{X}}^{\mathrm{price}}_{r_{\mathrm{in}}}\in\mathbb{R}^{(L+\mathcal{T})\times f_{1}},$ (1)
$\tilde{\mathbf{X}}^{\mathrm{price}}_{r_{\mathrm{in}}}$ $\xrightarrow{\mathrm{Dense}\ast k}\hat{\mathbf{X}}^{\mathrm{price}}_{r_{\mathrm{in}}}\in\mathbb{R}^{(L+\mathcal{T})\times h},$ (2)
${\mathbf{X}}^{\mathrm{exo}}_{r_{\mathrm{in}}}$ $\xrightarrow{\mathrm{Dense}\ast k}\hat{\mathbf{X}}^{\mathrm{exo}}_{r_{\mathrm{in}}}\in\mathbb{R}^{(L+\mathcal{T})\times h}.$ (3)
Next, we perform feature fusion through residual addition and flatten the fused representation to produce the regional representation $\mathbf{F}_{r_{\mathrm{in}}}$, which encodes the backward-looking price feature and forward-looking contextual information:
$\mathbf{F}_{r_{\mathrm{in}}}=\mathrm{Flatten}\left(\hat{\mathbf{X}}^{\mathrm{price}}_{r_{\mathrm{in}}}+\hat{\mathbf{X}}^{\mathrm{exo}}_{r_{\mathrm{in}}}\right)\in\mathbb{R}^{(L+\mathcal{T})\cdot h}.$ (4)
Next, the fused vectors $\mathbf{F}_{r_{\mathrm{in}}}$ for all regions $r_{\mathrm{in}}$$\in\mathcal{R}$ are stacked to form a spatial representation:
$\mathbf{S}=\mathrm{Stack}\left(\{\mathbf{F}_{r_{\mathrm{in}}}\}_{{r_{\mathrm{in}}}\in\mathcal{R}}\right)\in\mathbb{R}^{|\mathcal{R}|\times(L+\mathcal{T})\cdot h}.$ (5)
This fusion process enables feature aggregation across input regions with varying temporal and feature dimensions.
Graph Block
We first construct graph distance by performing a breadth-first search (BFS) traversal on the cross-border grid topology, which records the prior graph knowledge. For a given output region $r_{\mathrm{out}}\in\mathcal{R}$, we define the graph distance $d(r_{\mathrm{in}},r_{\mathrm{out}})$ as the minimal number of transmission hops from each input region $r_{\mathrm{in}}$ to out region $r_{\mathrm{out}}$, based on direct or indirect physical connectivity:
$d(r_{\mathrm{in}},r_{\mathrm{out}})=\begin{cases}0&\text{if }r_{\mathrm{in}}=r_{\mathrm{out}},\\ 1&\text{if }r_{\mathrm{in}}\sim r_{\mathrm{out}},\\ 1+\min\limits_{r^{\prime}\sim r_{\mathrm{in}}}d(r^{\prime},r_{\mathrm{out}})&\text{otherwise},\end{cases}$ (6)
where $r_{\mathrm{in}}\sim r_{\mathrm{out}}$ denotes that two regions are directly connected by a transmission line. For example, let $r_{\mathrm{out}}=\text{AT}$. Then $d(\text{AT},\text{AT})=0$. The region HU is directly connected to AT, thus $d(\text{HU},\text{AT})=1$. SK is not directly connected to AT but directly connected to HU, yielding $d(\text{SK},\text{AT})=2$.
Next, we introduce a graph decay mechanism to incorporate spatial prior knowledge. Intuitively, input regions that are topologically closer to the output region $r_{\mathrm{out}}$ are expected to exert a stronger influence than more distant ones. To formalize this intuition, we design a decay function that modulates the contribution of each neighboring region based on its graph distance, yielding a decay weight:
$w(d;c;D)\;=\;\begin{cases}\displaystyle\frac{(1-|c|)^{d}-(1-|c|)^{D}}{1-(1-|c|)^{D}},&c>0,\\ \displaystyle 1-\frac{d}{D},&c=0,\\ \displaystyle\frac{(1-|c|)^{-d}-(1-|c|)^{-D}}{1-(1-|c|)^{-D}},&c<0,\end{cases}$ (7)
where $w(d;c;D)\in[0,1]$ is the decay weight, $d$ is the graph distance, $c\in[-1,1]$ is the curvature parameter, and $D$ is the maximum reachable distance from the output region $r_{\mathrm{out}}$. As shown in Figure 4, for $c>0$ (Regime I), the decay weight drops off sharply with distance, meaning distant regions contribute very little. For $c=0$ (Regime II), the function reduces to linear decay, decreasing proportionally with graph distance. For $c<0$ (Regime III), the decay weight decreases more gradually, preserving the influence of neighbors. Using this decay function, the decay mask is constructed as:
$\mathbf{W}_{r_{\mathrm{out}}}=\begin{bmatrix}w(d(\text{AT},r_{\mathrm{out}});c;D)\\ w(d(\text{BE},r_{\mathrm{out}});c;D)\\ \vdots\\ w(d(\text{SK},r_{\mathrm{out}});c;D)\end{bmatrix}\in\mathbb{R}^{|\mathcal{R}|\times 1}.$ (8)
Next, the learned spatial representation $\mathbf{S}$ is copied $|\mathcal{R}|$ times, each assigned to an output region $r_{\mathrm{out}}$. We inject graph knowledge into $\mathbf{S}$ by computing the decay-weighted average representation:
$\mathbf{U}_{r_{\mathrm{out}}}=\frac{\mathbf{W}_{r_{\mathrm{out}}}^{\top}\mathbf{S}}{\mathbf{W}_{r_{\mathrm{out}}}^{\top}\mathbf{1}},$ (9)
where $\mathbf{W}_{r_{\mathrm{out}}}\in\mathbb{R}^{|\mathcal{R}|\times 1}$ is the decay mask vector for $r_{\mathrm{out}}$, and $\mathbf{1}\in\mathbb{R}^{|\mathcal{R}|\times 1}$ is a vector of ones.
This operation acts as spatial regularization and eliminates the need for an exhaustive learning process to determine spatial weights, as required in methods such as attention.
Head
We design a multi-region, multi-timestep, and multi-quantile head, where the model produces joint probabilistic forecasts. To prevent quantile crossing issue [^1] Quantile crossing refers to the phenomenon where upper quantile predictions (e.g., 90%) fall below lower quantiles (e.g., 10%), violating the monotonicity of the quantile function. (Chernozhukov, Fernández-Val, and Galichon 2010)., we adopt a hierarchical quantile head (Yu et al. 2025). In detail, the median quantile ($\tau=0.5$) price trajectory, which represents the full set of timesteps $\mathcal{T}$, is predicted from $\mathbf{U}_{r_{\mathrm{out}}}$ via a dense layer:
$\hat{\mathbf{y}}_{{r_{\mathrm{out}}},0.5}=\text{Dense}(\mathbf{U}_{r_{\mathrm{out}}})\in\mathbb{R}^{\mathcal{T}}.$ (10)
For the upper quantile ($\tau=0.9$), a residual price trajectory $\hat{\mathbf{r}}_{{r_{\mathrm{out}}},0.9}$ is predicted from $\mathbf{U}_{r_{\mathrm{out}}}$ via a dense layer:
$\hat{\mathbf{r}}_{{r_{\mathrm{out}}},0.9}=\text{Dense}(\mathbf{U}_{r_{\mathrm{out}}})\in\mathbb{R}^{\mathcal{T}},$ (11)
where a non-negative function $g(\cdot)$, such as absolute-value function, is applied to the price residual. The upper quantile forecast is obtained by adding this non-negative residual to the median prediction:
$\hat{\mathbf{y}}_{{r_{\mathrm{out}}},0.9}=\hat{\mathbf{y}}_{{r_{\mathrm{out}}},0.5}+g(\hat{\mathbf{r}}_{{r_{\mathrm{out}}},0.9}).$ (12)
For the lower quantile ($\tau=0.1$), we compute a residual trajectory similarly:
$\hat{\mathbf{r}}_{{r_{\mathrm{out}}},0.1}=\text{Dense}(\mathbf{U}_{r_{\mathrm{out}}})\in\mathbb{R}^{\mathcal{T}},$ (13)
and subtract it from the median to obtain the lower quantile prediction:
$\hat{\mathbf{y}}_{{r_{\mathrm{out}}},0.1}=\hat{\mathbf{y}}_{{r_{\mathrm{out}}},0.5}-g(\hat{\mathbf{r}}_{{r_{\mathrm{out}}},0.1}).$ (14)
This hierarchical design guarantees that the upper quantile prediction is greater than or equal to the lower one at each time step, effectively overcoming quantile crossing.
Loss
We introduce the Average Quantile Loss (AQL) as the training objective for multi-region, multi-timestep, and multi-quantile probabilistic forecasting. Let $y_{i,{r_{\mathrm{out}}},t}$ denote the ground-truth price for the $i$-th training sample, output region $r_{\mathrm{out}}$}, and timestep $t$, and let $\hat{y}_{i,{r_{\mathrm{out}}},t,\tau}$ be the corresponding predicted quantile. The AQL across all training samples, regions, timesteps, and quantiles is computed as:
$\mathrm{AQL}=\frac{1}{N\left|\mathcal{R}\right|\mathcal{T}\left|\mathcal{Q}\right|}\sum_{i=1}^{N}\sum_{r_{\mathrm{out}}\in\mathcal{R}}\sum_{t=1}^{\mathcal{T}}\sum_{\tau\in\mathcal{Q}}L_{\tau}\left(y_{i,r_{\mathrm{out}},t},\hat{y}_{i,r_{\mathrm{out}},t,\tau}\right),$ (15)
where $N$ is the number of samples, and the quantile loss function $L_{\tau}$ for a single prediction task is defined as:
$L_{\tau}(y,\hat{y}_{\tau})=\begin{cases}\tau\cdot(y-\hat{y}_{\tau}),&\text{if }y\geq\hat{y}_{\tau},\\ (1-\tau)\cdot(\hat{y}_{\tau}-y),&\text{otherwise},\end{cases}$ (16)
where $y$ and $\hat{y}$ are the true and predicted values, respectively.
A lower AQL indicates more accurate probabilistic forecasting across spatial and temporal dimensions.
Other Details
We use the Adam optimizer with an initial learning rate of $4\times 10^{-3}$, which decays exponentially by a factor of 0.95 every 10 epochs. The batch size is 8, which introduces a slight regularization effect. The model is trained for 50 epochs, and the checkpoint with the lowest validation loss is saved.
Experiment
For the three-year dataset, we use the first two years of the dataset (67%) for training, and the final year (33%) is evenly split into validation and testing sets, strictly preserving the temporal order to ensure realistic out-of-sample evaluation. We assess the model performance using the quantile loss at individual levels (Q10, Q50, and Q90), AQL, Average Quantile Crossing Rate (AQCR), Root Mean Squared Error (RMSE), Mean Absolute Error (MAE), and coefficient of determination ($R^{2}$). The Diebold-Mariano (DM) test is applied to determine if two models have a significant difference in the probabilistic and pointwise forecasting (Diebold and Mariano 2002). Each metric is explained in Appendix, Section Metrics.
Model Comparison
We include several spatial and temporal models as baselines. Model descriptions and hyperparameter optimization are detailed in Appendix, Section Baselines.
Spatial Models
Following (Yang et al. 2024; Meng et al. 2024), where variants of GNNs, such as the Graph Convolutional Network (GraphConv) and the Graph Attention Network (GraphAttn), have made successes in electricity price forecasting, we include them as baselines. We also compare with other advanced variants, such as a GNN based on sample-and-aggregate operations (GraphSAGE) (Hamilton, Ying, and Leskovec 2017), a diffusion-based GNN (GraphDiff) (Li et al. 2018), and a GNN with Auto-Regressive Moving Average filters (GraphARMA) (Bianchi et al. 2021).
Results in Table 1 show that PriceFM statistically outperforms all spatial models, confirmed by both the probabilistic and pointwise DM test. A common limitation of these models is that they have only limited spatial priors, relying primarily on data-driven learning to assign spatial importance. Without an explicit mechanism to enforce spatial decay, they lack regularization over noisy regions. As a result, they require large datasets to generalize. However, despite our dataset spanning three years, which is considered large in the domain, the daily forecasting requirement limits the training set to only around 700 samples, making it unsuitable for such models.
Model Q10 $\downarrow$ Q50 $\downarrow$ Q90 $\downarrow$ AQL $\downarrow$ AQCR $\downarrow$ RMSE $\downarrow$ MAE $\downarrow$ ${R}^{2}$ $\uparrow$
GraphConv 4.81±0.16 10.76±0.16 7.11±0.33 7.56±0.15 2.19±0.76 35.75±0.60 21.53±0.33 0.51±0.02
GraphAttn 5.08±0.31 11.58±0.37 7.90±0.40 8.18±0.29 1.41±0.47 37.63±0.62 23.15±0.75 0.44±0.01
GraphSAGE 5.20±0.13 11.27±0.23 7.39±0.35 7.95±0.18 2.79±0.56 37.09±0.81 22.53±0.46 0.47±0.02
GraphDiff 4.80±0.20 11.03±0.21 7.34±0.33 7.73±0.18 3.33±0.70 36.35±0.80 22.07±0.42 0.48±0.02
GraphARMA 4.87±0.07 11.10±0.22 7.00±0.22 7.66±0.16 2.05±0.65 36.15±0.59 22.21±0.44 0.49±0.02
FEDFormer 47.55±1.51 28.91±0.99 0.14±0.00
TimesNet 44.20±0.87 26.40±0.52 0.30±0.01
TimeXer 47.38±1.18 28.29±0.71 0.14±0.02
PriceFM 4.89±0.05 9.83±0.15 5.96±0.08 6.89±0.08 0.00±0.00 32.25±0.38 19.65±0.30 0.61±0.01
Table 1: Model performance comparison on the testing set. The symbol “–” indicates that the model does not support probabilistic forecasting by design. All metrics are reported as mean±standard deviation over 5 independent runs. The best results are shown in bold, and the second-best are underlined. The units of Q10, Q50, Q90, AQL, RMSE, and MAE are expressed in €/MWh (denormalized), and AQCR in %.
Temporal Models
We include FEDFormer (Zhou et al. 2022), TimesNet (Wu et al. 2023), and TimeXer (Wang et al. 2024) in our evaluation. As these foundation models do not support graph-based input, features from all regions are concatenated along the feature dimension. Including these models aims to investigate whether these pure time-series models can capture spatial patterns without prior graph knowledge.
Results in Table 1 show that PriceFM statistically outperforms all temporal models, confirmed by the pointwise DM test. In particular, the inclusion of noisy features from weakly related regions easily leads to overfitting, given the high model complexity of these time-series foundation models. This observation confirms that the temporal models struggle to recognize spatial patterns without explicit injection of graph knowledge. In contrast, PriceFM incorporates a graph decay mechanism that acts as a spatial regularizer, thereby attenuating the influence of noisy regions.
Ablation Study
We systematically evaluate the impact of key hyperparameters and architectural design choices, including spatiotemporal configurations, fusion strategies in the fusion block, graph decay mechanisms in the graph block, and variants of the hierarchical quantile head.
Spatiotemporal Configurations
We assess the impact of two key hyperparameters that control the spatial decay strength and temporal input length:
  • Curvature Parameter: Spatially, we evaluate $c\in\{-1.0,-0.8,\ldots,0.8,1.0\}$ in increments of 0.2, ranging from weak decay to strong decay. In total, 2,090 trials are conducted to determine the optimal curvature value for each output region individually, based on validation loss.
  • Backward-Looking Window Size: Temporally, we compare $L\in\{24,72,168\}$, corresponding to one day, three days, and one week. For each window size, all other hyperparameters are re-optimized.
Figure 5 illustrates the testing loss and the distribution of optimal curvature values across all regions. Most regions confirm spatial interdependencies ($c\neq 1.0$). The identified optimal curvature shows a weak negative correlation ($\rho=-0.18$) with the number of directly connected regions, mentioned in Figure 1. The results in Table 2 indicate that the optimal backward-looking window size is 24, potentially because information from the distant past becomes outdated.
Figure 5: Spatial distribution of testing loss and curvature values. a Average quantile loss per region on the testing set. Western European regions exhibit lower losses, whereas BG, HU, and RO show particularly high losses (orange areas). b Optimal curvature value per region. Particularly, 8 regions from IT and SE have a curvature value of 1, indicating optimal performance by excluding neighboring features.
Method Q10 $\downarrow$ Q50 $\downarrow$ Q90 $\downarrow$ AQL $\downarrow$ AQCR $\downarrow$ RMSE $\downarrow$ MAE $\downarrow$ ${R}^{2}$ $\uparrow$
$L=24^{\dagger}$ 4.89±0.05 9.83±0.15 5.96±0.08 6.89±0.08 0.00±0.00 32.25±0.38 19.65±0.30 0.61±0.01
$L=72$ 5.09±0.13 10.04±0.20 6.27±0.06 7.14±0.09 0.00±0.00 32.76±0.41 20.09±0.39 0.60±0.01
$L=168$ 5.62±0.30 10.52±0.26 6.65±0.16 7.60±0.21 0.00±0.00 33.82±0.59 21.04±0.53 0.57±0.01
Res. Add 4.89±0.05 9.83±0.15 5.96±0.08 6.89±0.08 0.00±0.00 32.25±0.38 19.65±0.30 0.61±0.01
Concat. 5.03±0.18 10.50±0.21 6.40±0.06 7.31±0.13 0.00±0.00 34.68±0.49 21.01±0.42 0.56±0.01
Cross-Attn 4.92±0.14 9.76±0.10 6.07±0.09 6.92±0.09 0.00±0.00 32.46±0.27 19.51±0.19 0.61±0.01
Decay 4.89±0.05 9.83±0.15 5.96±0.08 6.89±0.08 0.00±0.00 32.25±0.38 19.65±0.30 0.61±0.01
Random 5.32±0.27 10.95±0.13 7.21±0.11 7.83±0.11 0.00±0.00 35.85±0.30 21.89±0.25 0.50±0.02
No Decay 5.23±0.32 11.05±0.22 6.84±0.22 7.71±0.16 0.00±0.00 35.81±0.63 22.11±0.44 0.50±0.02
ABS 4.89±0.05 9.83±0.15 5.96±0.08 6.89±0.08 0.00±0.00 32.25±0.38 19.65±0.30 0.61±0.01
Square 4.86±0.14 9.80±0.16 6.05±0.21 6.90±0.15 0.00±0.00 32.35±0.36 19.59±0.33 0.61±0.01
ReLU 4.76±0.15 9.84±0.07 6.06±0.14 6.89±0.07 0.00±0.00 32.48±0.13 19.68±0.15 0.61±0.00
Standard 4.80±0.14 10.05±0.06 5.94±0.13 6.93±0.04 4.10±1.28 32.76±0.18 20.10±0.13 0.60±0.01
Table 2: Ablation studies of different module choices: temporal configuration (rows 1-3), fusion block (rows 4–6), graph block (rows 7–9), and hierarchical quantile head (rows 10–13). The symbol marks the method used in PriceFM. All metrics are reported as mean±standard deviation over 5 independent runs. The best results are shown in bold, and the second-best are underlined. The units of Q10, Q50, Q90, AQL, RMSE, and MAE are expressed in €/MWh (denormalized), and AQCR in %.
Fusion Block
We compare the residual addition from the fusion block, described in Eq. (4), with two fusion strategies:
  • Concatenation: We replace Eq. (4) by first flattening both $\hat{\mathbf{X}}^{\mathrm{price}}_{r_{\mathrm{in}}}$ and $\hat{\mathbf{X}}^{\mathrm{exo}}_{r_{\mathrm{in}}}$, then concatenating them:
    $\mathbf{F}_{r_{\mathrm{in}}}=\mathrm{Concat}\left(\mathrm{Flatten}\left(\hat{\mathbf{X}}^{\mathrm{price}}_{r_{\mathrm{in}}}\right),\,\mathrm{Flatten}\left(\hat{\mathbf{X}}^{\mathrm{exo}}_{r_{\mathrm{in}}}\right)\right).$ (17)
  • Cross-Attention: We apply multi-head attention with $\hat{\mathbf{X}}^{\mathrm{price}}_{r_{\mathrm{in}}}$ as the query and $\hat{\mathbf{X}}^{\mathrm{exo}}_{r_{\mathrm{in}}}$ as both key and value to produce $\hat{\mathbf{X}}^{\mathrm{attn}}_{r_{\mathrm{in}}}$. The attended features are then fused back into the price representation using residual addition:
    $\mathbf{F}_{r_{\mathrm{in}}}=\mathrm{Flatten}\left(\hat{\mathbf{X}}^{\mathrm{price}}_{r_{\mathrm{in}}}+\hat{\mathbf{X}}^{\mathrm{attn}}_{r_{\mathrm{in}}}\right).$ (18)
The results in Table 2 show that replacing the residual addition with concatenation leads to 6.10% higher AQL for probabilistic prediction. Switching to cross-attention yields comparable performance to residual addition, while introducing additional model parameters and unnecessary complexity. This suggests that the residual addition strikes a favorable balance between performance and model simplicity.
Graph Block
We compare the graph decay mask from the graph block, described in Eq. (8), with two alternatives:
  • Random Graph Decay Mask: We replace Eq. (8) with a randomly sampled vector, where each decay weight is drawn independently from a uniform distribution over $[0,1]$, thereby removing the spatial graph prior:
    $\mathbf{W}_{r_{\mathrm{out}}}\sim\mathcal{U}(0,1)^{|\mathcal{R}|\times 1}.$ (19)
  • No Graph Decay: We remove the decay mask, which simplifies Eq. (9) to a uniform average over input regions:
    $\mathbf{U}_{r_{\mathrm{out}}}=\frac{\mathbf{1}^{\top}\mathbf{S}}{|\mathcal{R}|},$ (20)
The results in Table 2 demonstrate that randomizing or removing the graph decay mask, which discards the prior graph knowledge, leads to a significant drop in all metrics. We also observe that such results, obtained without prior graph knowledge, are on par with those of GNN variants, shown in Table 1. We emphasize that relying on pure data-driven learning without an explicit decay mechanism leads to a loss of the key inductive bias, limiting the model’s performance, especially when the training data is scarce.
Hierarchical Quantile Head
We evaluate the impact of different non-negative functions used in Eq. (12) and (14), and include a standard multi-quantile head for comparison.
  • Non-Negative Functions: We replace the absolute-value function with either a square function or ReLU:
    $g(\cdot)=(\cdot)^{2},$ (21)
    $g(\cdot)=\max(0,\cdot).$ (22)
  • Standard Multi-Quantile Head: The Eq. (11) and (13) are skipped, and the shared representation $\mathbf{U}_{r_{\mathrm{out}}}$ is passed directly to three independent dense layers to produce each quantile trajectory:
    $\hat{\mathbf{y}}_{r_{\mathrm{out}},0.1}=\mathrm{Dense}(\mathbf{U}_{r_{\mathrm{out}}})\in\mathbb{R}^{\mathcal{T}},$ (23)
    $\hat{\mathbf{y}}_{r_{\mathrm{out}},0.5}=\mathrm{Dense}(\mathbf{U}_{r_{\mathrm{out}}})\in\mathbb{R}^{\mathcal{T}},$ (24)
    $\hat{\mathbf{y}}_{r_{\mathrm{out}},0.9}=\mathrm{Dense}(\mathbf{U}_{r_{\mathrm{out}}})\in\mathbb{R}^{\mathcal{T}}.$ (25)
The results in Table 2 reveal that replacing the absolute-value function with either a square function or ReLU does not result in a noticeable change in performance, suggesting that the choice of non-negative function is flexible. Moreover, while the hierarchical quantile head achieves comparable loss to the standard multi-quantile head, the latter exhibits a mean AQCR of 4.10%, indicating that the hierarchical design mitigates quantile crossing without harming performance.
Conclusion
In this paper, we introduced and released a dataset, which will benefit both the research community and the energy industry. We proposed and released PriceFM, a foundation model tailored for interconnected European electricity markets. Extensive experiments and ablation studies demonstrate that PriceFM outperforms competitive baselines and highlight the importance of spatial context. By enabling more accurate and comprehensive probabilistic electricity price forecasting, our work has the potential to support better decision-making in energy trading and grid management, contributing to a more efficient and resilient energy ecosystem.
Despite these promising results, our study has two key limitations. First, the graph decay function is empirically designed; the performance may improve with alternative functions. Second, PriceFM relies on the current transmission topology, which may become outdated as transmission networks evolve, requiring retraining to adapt to new scenarios.
Appendix
Code Guideline
We open-source all code for preprocessing, modeling, and analysis. The project directory is structured as follows:
|- PriceFM/
   |- Data/
   |- Model/
   |- Result/
   |- PriceFM.py
   |- Main.py
   |- README.md
where the README.md specifies the required package version. To facilitate reproducibility and accessibility, we have streamlined the entire pipeline through extensive engineering efforts into just three simple steps:
Step 1: Create a folder named PriceFM, along with subfolders Data, Model, and Result. Place the energy data EU_Spatiotemporal_Energy_Data.csv into Data, and place PriceFM.py inside the PriceFM folder.
Step 2: Run Main.py to process the energy data, and to train, validate, and test the PriceFM. The script PriceFM.py contains all necessary functions and classes.
Step 3: After execution, check the Model and Result folders for the saved model weights and evaluation metrics.
Hardware and Computation
The PriceFM is evaluated on both an NVIDIA A100 GPU and an Intel Core i7-1265U CPU, respectively. The NVIDIA A100 is designed for high-performance computing and deep learning workloads, offering 80 GB of high-bandwidth memory and up to 6,912 CUDA cores. In contrast, the Intel i7-1265U is a power-efficient CPU commonly found in standard laptops. This dual evaluation setup allows us to assess the feasibility of deploying our model in environments without access to advanced GPU hardware.
Under the training setup described in Section Model, the training time is approximately 4–5 minutes on the A100 GPU and 12–13 minutes on the i7 CPU. Due to the small batch size used, the GPU does not achieve a significant speedup over the CPU. Inference time for both setups is under 10 seconds. We note that neither training nor inference time is critical for our application, as bid submissions can occur at any point before the market gate closure.
Lookup Table and Feature Availability
The country-region code lookup table and the feature availability are listed in Table 1.
Country Region Code Price Load Solar Wind (Onshore) Wind (Offshore)
Austria AT
Belgium BE
Bulgaria BG
Czech Republic CZ
Germany, Luxembourg DE-LU
Denmark DK1
Denmark DK2
Estonia $E$
Spain ES
Finland FI
France FR (✓)
Greece GR
Croatia HR
Hungary HU
Italy IT-CALA
Italy IT-CNOR
Italy IT-CSUD
Italy IT-NORD
Italy IT-SARD
Italy IT-SICI
Italy IT-SUD
Lithuania LT
Latvia LV (✓)
Netherlands NL
Norway NO1
Norway NO2
Norway NO3
Norway NO4
Norway NO5
Poland PL
Portugal PT
Romania RO
Sweden SE1
Sweden SE2
Sweden SE3
Sweden SE4
Slovenia SI
Slovakia SK (✓)
Table 1: Lookup table and feature availability across European regions. ✓ indicates that the feature is available. (✓) denotes partial availability, and the feature is excluded from this study due to the high rate of missing values.
Baselines
We compare the PriceFM with several spatial and temporal models. The PriceFM is optimized based on validation loss, and the hyperparameter search space is summarized in Table 2. The number of parameters of PriceFM is only 3.38M.
The spatial models, such as GraphConv, GraphAttn, GraphDiff, GraphSAGE, and GraphARMA, lack an explicit spatial decay mechanism guided by graph distance. GraphConv assigns uniform weights to all neighbors, ignoring spatial relevance; GraphAttn learns attention weights entirely from data without structural priors; GraphDiff uses diffusion kernels that spread information globally, but the importance of nodes is still determined through learned weights, without decay constraints; GraphSAGE aggregates features from sampled neighbors but lacks a notion of spatial proximity; and GraphARMA applies recursive smoothing, which can propagate noise from irrelevant neighbors. As a result, these models require large datasets to recognize spatial patterns and may struggle to suppress the influence of noisy regions. The hyperparameters are optimized based on validation loss, and the search space is summarized in Table 3. All spatial models require an adjacency matrix as input, detailed in Section Spatial Topology and Adjacency Matrix. Notably, these optimized spatial models contain more than 10M parameters.
FEDFormer, TimesNet, and TimeXer have demonstrated strong performance in general time-series forecasting tasks. FEDFormer and TimeXer are Transformer-based architectures, while TimesNet is CNN-based. However, these models are not designed for spatial graph forecasting tasks and may require large-scale data to implicitly learn spatial dependencies. The hyperparameters are optimized based on validation loss, and the search space is summarized in Table 4. Notably, the number of parameters of these optimized temporal models ranges from 5M to 10M, and they are limited to single-region prediction. As a result, 38 separate models must be trained for each output region, since they are not designed to support multi-region forecasting.
Table 2: Hyperparameter search space for PriceFM.
Model Search Space
PriceFM
hidden_size: {12, 24, 48}
layers: {2, 3, 4}
batch_size: {8, 32, 128}
learning_rate: {4e-4, 1e-3, 4e-3}
epochs: 50
Table 3: Hyperparameter search space for spatial models.
Model Search Space
GraphConv
hidden_size: {32, 128, 512}
layers: {2, 3, 4}
batch_size: {8, 32, 128}
learning_rate: {4e-4, 1e-3, 4e-3}
epochs: 50
GraphAttn
hidden_size: {32, 128, 512}
layers: {2, 3, 4}
n_heads: {2, 4, 8}
batch_size: {8, 32, 128}
dropout: {0.1, 0.3, 0.5}
learning_rate: {4e-4, 1e-3, 4e-3}
epochs: 50
GraphSAGE
hidden_size: {32, 128, 512}
layers: {2, 3, 4}
aggregate: {mean, max, sum}
batch_size: {8, 32, 128}
learning_rate: {4e-4, 1e-3, 4e-3}
epochs: 50
GraphDiff
diff_steps: {2, 4, 6}
hidden_size: {32, 128, 512}
layers: {2, 3, 4}
batch_size: {8, 32, 128}
learning_rate: {4e-4, 1e-3, 4e-3}
epochs: 50
GraphARMA
hidden_size: {32, 128, 512}
layers: {2, 3, 4}
order: {1, 2, 4}
iteration: {1, 2, 4}
batch_size: {8, 32, 128}
learning_rate: {4e-4, 1e-3, 4e-3}
epochs: 50
Table 4: Hyperparameter search space for temporal models.
Model Search Space
FEDFormer
hidden_size: {32, 128, 512}
conv_hidden_size: {32, 128, 512}
e_layers: {2, 3, 4}
n_heads: {2, 4, 8}
dropout: {0.1, 0.3, 0.5}
batch_size: {8, 32, 128}
learning_rate: {4e-4, 1e-3, 4e-3}
epochs: 50
TimesNet
hidden_size: {32, 128, 512}
conv_hidden_size: {32, 128, 512}
e_layers: {2, 3, 4}
dropout: {0.1, 0.3, 0.5}
batch_size: {8, 32, 128}
learning_rate: {4e-4, 1e-3, 4e-3}
Epochs: 50
TimeXer
hidden_size: {32, 128, 512}
e_layers: {2, 3, 4}
n_heads: {2, 4, 8}
d_ff: {512, 1024, 2048}
dropout: {0.1, 0.3, 0.5}
batch_size: {8, 32, 128}
learning_rate: {4e-4, 1e-3, 4e-3}
epochs: 50
Spatial Topology and Adjacency Matrix
We model the European market as a graph $G=(\mathcal{R},\mathcal{E})$, where each node $r\in\mathcal{R}$ is a bidding zone and edges indicate direct power flow via cross-border interconnections. This spatial topology is detailed in Table 5.
Let $\mathcal{N}(r)$ denote the set of directly connected neighbors of $r$, excluding $r$ itself. The binary adjacency matrix $A\in\{0,1\}^{|\mathcal{R}|\times|\mathcal{R}|}$ is defined by
$A_{r,s}=\begin{cases}1,&\text{if }s\in\mathcal{N}(r),\\ 0,&\text{otherwise,}\end{cases}\quad r,s\in\mathcal{R}.$ (26)
For GNN layers, self-loops can be added via $\tilde{A}=A+I$.
Scaling
To normalize the data while being robust to extreme values, we employ a RobustScaler fitted on the training data, using the Scikit-Learn implementation. The fitted scaler is then used to transform validation and testing data.
Metrics
We evaluate the model using Q10, Q50, Q90, AQL, AQCR, RMSE, MAE, and $R^{2}$, all averaged over regions and samples.
Q10, Q50, Q90
We compute quantile loss separately for each target quantile. For example, the Q10 loss is:
$\mathrm{Q10}=\frac{1}{N\left|\mathcal{R}\right|\mathcal{T}}\sum_{i=1}^{N}\sum_{r_{\mathrm{out}}\in\mathcal{R}}\sum_{t=1}^{\mathcal{T}}L_{0.1}\left(y_{i,r_{\mathrm{out}},t},\hat{y}_{i,r_{\mathrm{out}},t,0.1}\right),$ (27)
and similarly for Q50 and Q90 with $\tau=0.5$ and $0.9$.
AQL
AQL represents the average quantile loss across all quantiles, as described in Section Model.
AQCR
AQCR captures the proportion of forecasted distributions that violate quantile monotonicity, i.e., when a higher quantile is predicted to be smaller than a lower one. The quantile crossing indicator for a quantile pair $(\tau_{l},\tau_{u})$ with $\tau_{l}<\tau_{u}$ is defined as:
$C_{\tau_{l},\tau_{u}}(i,r,t)=\mathbb{1}\left[\hat{y}_{i,r,t,\tau_{l}}>\hat{y}_{i,r,t,\tau_{u}}\right],$ (28)
where $\mathbb{1}[\cdot]$ is the indicator function. Then, the AQCR is given by:
$\mathrm{AQCR}=\frac{1}{N\left|\mathcal{R}\right|\mathcal{T}}\sum_{i=1}^{N}\sum_{r\in\mathcal{R}}\sum_{t=1}^{\mathcal{T}}\mathrm{Cross}_{i,r,t}.$ (29)
RMSE
We compute RMSE within each region, then average over all regions:
$\mathrm{RMSE}_{r}=\sqrt{\frac{1}{N\mathcal{T}}\sum_{i=1}^{N}\sum_{t=1}^{\mathcal{T}}\left(y_{i,r,t}-\hat{y}_{i,r,t,0.5}\right)^{2}},$ (30)
$\mathrm{RMSE}=\frac{1}{\left|\mathcal{R}\right|}\sum_{r\in\mathcal{R}}\mathrm{RMSE}_{r}.$ (31)
MAE
Same procedure as RMSE, but using absolute error:
$\mathrm{MAE}_{r}=\frac{1}{N\mathcal{T}}\sum_{i=1}^{N}\sum_{t=1}^{\mathcal{T}}\left|y_{i,r,t}-\hat{y}_{i,r,t,0.5}\right|,$ (32)
$\mathrm{MAE}=\frac{1}{\left|\mathcal{R}\right|}\sum_{r\in\mathcal{R}}\mathrm{MAE}_{r}.$ (33)
$R^{2}$
We compute $R^{2}$ for each region and average across all regions:
$R^{2}_{r}=1-\frac{\sum_{i=1}^{N}\sum_{t=1}^{\mathcal{T}}\left(y_{i,r,t}-\hat{y}_{i,r,t,0.5}\right)^{2}}{\sum_{i=1}^{N}\sum_{t=1}^{\mathcal{T}}\left(y_{i,r,t}-\bar{y}_{r}\right)^{2}},$ (34)
$\bar{y}_{r}=\frac{1}{N\mathcal{T}}\sum_{i=1}^{N}\sum_{t=1}^{\mathcal{T}}y_{i,r,t},$ (35)
$R^{2}=\frac{1}{\left|\mathcal{R}\right|}\sum_{r\in\mathcal{R}}R^{2}_{r}.$ (36)
DM Test
To assess whether differences in forecasting performance are statistically significant, we apply the DM test. We compute the loss differential at each prediction instance.
For probabilistic forecasts, we compute the loss differential at each quantile $\tau\in\mathcal{Q}$ between two models $l\in\{1,2\}$:
$d_{i,r,t,\tau}=L_{\tau}\left(y_{i,r,t},\hat{y}^{(1)}_{i,r,t,\tau}\right)-L_{\tau}\left(y_{i,r,t},\hat{y}^{(2)}_{i,r,t,\tau}\right)$. (37)
For point forecasts, the loss differential between two models is computed for each sample, region, and timestep as:
[…]
The DM test statistic is then calculated as:
$\mathrm{DM}=\frac{\bar{d}}{\hat{\sigma}_{d}/\sqrt{M}},$ (39)
$\bar{d}=\frac{1}{M}\sum_{j=1}^{M}d_{j},$ (40)
where $M=N\cdot|\mathcal{R}|\cdot\mathcal{T}\cdot|\mathcal{Q}|$ for probabilistic forecasts, and $M=N\cdot|\mathcal{R}|\cdot\mathcal{T}$ for point forecasts. The index $j$ enumerates all prediction instances across dimensions, and $\hat{\sigma}_{d}$ is the sample standard deviation of $\{d_{j}\}_{j=1}^{M}$.
We compute a two-sided $p$-value using the Student’s $t$-distribution with M1M-1 degrees of freedom. A significantly positive (negative) DM value indicates that model 2 (model 1) significantly outperforms the other.
Table 5: Direct neighbors by region and neighbor count.
Region Code Direct Neighbors Count
AT CZ, DE-LU, HU, IT-NORD, SI 5
BE DE-LU, FR, NL 3
BG GR, RO 2
CZ AT, DE-LU, PL, SK 4
DE-LU AT, BE, CZ, DK1, DK2, FR, NL, NO2, PL, SE4 10
DK1 DE-LU, DK2, NL, NO2, SE3 5
DK2 DE-LU, DK1, SE4 3
$E$ FI, LV 2
ES FR, PT 2
FI EE, NO4, SE1, SE3 4
FR BE, DE-LU, ES, IT-NORD 4
GR BG, IT-SUD 2
HR HU, SI 2
HU AT, HR, RO, SI, SK 5
IT-CALA IT-SICI, IT-SUD 2
IT-CNOR IT-CSUD, IT-NORD 2
IT-CSUD IT-CNOR, IT-SARD, IT-SUD 3
IT-NORD AT, FR, IT-CNOR, SI 4
IT-SARD IT-CSUD 1
IT-SICI IT-CALA 1
IT-SUD GR, IT-CALA, IT-CSUD 3
LT LV, PL, SE4 3
LV EE, LT 2
NL BE, DK1, DE-LU, NO2 4
NO1 NO2, NO3, NO5, SE3 4
NO2 DE-LU, DK1, NL, NO1, NO5 5
NO3 NO1, NO4, NO5, SE2 4
NO4 FI, NO3, SE1, SE2 4
NO5 NO1, NO2, NO3 3
PL CZ, DE-LU, LT, SE4, SK 5
PT ES 1
RO BG, HU 2
SE1 FI, NO4, SE2 3
SE2 NO3, NO4, SE1, SE3 4
SE3 DK1, FI, NO1, SE2, SE4 5
SE4 DE-LU, DK2, LT, PL, SE3 5
SI AT, HR, HU, IT-NORD 4
SK CZ, HU, PL 3

### v2

@@
 Extensive experiments and ablation studies confirm the model’s effectiveness, consistently outperforming competitive baselines and highlighting the importance of spatial context in electricity markets.
-The dataset and code can be found at https://github.com/runyao-yu/PriceFM.
-Introduction
-The European electricity market is physically interconnected through a network of cross-border transmission lines, enabling the exchange of electricity between regions and optimizing the social welfare at the European level (Lago et al.
-2018).
-However, physical constraints, such as limited transmission capacity, can restrict electricity flow between regions and lead to zonal price differences (Finck 2021), illustrated in Figure 1.
+The project page is at: https://runyao-yu.github.io/PriceFM/.
+1 Introduction
+The European electricity market is physically interconnected through a network of cross-border transmission lines, enabling the exchange of electricity between regions and optimizing the social welfare at the European level Lago et al.
+(2018).
+However, physical constraints, such as limited transmission capacity, can restrict electricity flow between regions and lead to zonal price differences Finck (2021), illustrated in Figure 1.
 These price disparities highlight the spatial nature of electricity price formation.
-Recent studies show that electricity price dynamics are strongly influenced by spatial interdependencies and cannot be accurately captured using region-specific models (Do et al.
-2024).
+Recent studies show that electricity price dynamics are strongly influenced by spatial interdependencies and cannot be accurately captured using region-specific models Do et al.
+(2024).
 Therefore, explicitly modeling the spatial structure of the European electricity market is essential for producing accurate price forecasts.
 Figure 1: Spatial distribution of electricity price and number of neighboring regions.
-a Electricity prices for 38 European regions averaged from 2022-01-01 to 2025-01-01.
+(a) Electricity prices for 38 European regions averaged from 2022-01-01 to 2025-01-01.
 A significant zonal price difference is observed between north and south regions.
-b Number of neighboring regions that are directly connected to certain region via transmission lines.
-For example, France (FR) and Portugal (PT) are directly connected to Spain (ES), thus the number of regions for ES is 2.
-Most existing studies on electricity price forecasting do not explicitly model the spatial structure and focus on a single-region market, particularly Germany (Muniain and Ziel 2020; Maciejowska, Nitka, and Weron 2021; Kitsatoglou et al.
-2024), as the German market is one of the largest markets in Europe.
-Other studies explore forecasting methods for markets such as Denmark, Finland, and Spain, also using region-specific models (Ziel and Weron 2018; Gianfreda, Ravazzolo, and Rossini 2020; Loizidis, Kyprianou, and Georghiou 2024).
+(b) Number of neighboring regions that are directly connected to certain region via transmission lines.
+For example, France (FR) and Portugal (PT) are directly connected to Spain (ES), thus the number of neighboring regions for ES is 2.
+The mean value across all regions is 3.4.
+Most existing studies on electricity price forecasting do not explicitly model the spatial structure and focus on a single-region market, particularly Germany Muniain & Ziel (2020); Maciejowska et al.
+(2021); Kitsatoglou et al.
+(2024), as the German market is one of the largest markets in Europe.
+Other studies explore forecasting methods for markets such as Denmark, Finland, and Spain, also using region-specific models Ziel & Weron (2018); Gianfreda et al.
+(2020); Loizidis et al.
+(2024).
 More recent works explicitly model the spatial nature of the electricity price.
-For instance, a Graph Convolutional Network (GCN) is applied to capture spatial interdependencies in the Nordic markets, such as Norway, Sweden, and Finland (Yang et al.
-2024).
-Moreover, an attention-based variant is developed to predict prices in certain European markets such as Austria, Germany, and Hungary (Meng et al.
-2024).
+For instance, a Graph Convolutional Network (GCN) is applied to capture spatial interdependencies in the Nordic markets, such as Norway, Sweden, and Finland Yang et al.
+(2024).
+Moreover, an attention-based variant is developed to predict prices in certain European markets such as Austria, Germany, and Hungary Meng et al.
+(2024).
 However, these models cover only subsets of Europe and primarily produce point forecasts, failing to capture the uncertainty inherent in electricity prices.
-Uncertainty modeling in electricity markets is critical, as the electricity price is strongly influenced by intermittent renewable generation and fluctuating demand (Lago et al.
-2021).
-Consequently, price forecasting should extend beyond traditional pointwise forecasting to explicitly quantify uncertainties, especially for applications involving risk-sensitive applications such as energy trading and operational planning (Ziel and Steinert 2018).
-A comprehensive survey summarizes various probabilistic forecasting approaches based on quantile regression methods (Lago et al.
-2021).
+Uncertainty modeling in electricity markets is critical, as the electricity price is strongly influenced by intermittent renewable generation and fluctuating demand Lago et al.
+(2021).
+Consequently, price forecasting should extend beyond traditional pointwise forecasting to explicitly quantify uncertainties, especially for applications involving risk-sensitive applications such as energy trading and operational planning Ziel & Steinert (2018).
+A comprehensive survey summarizes various probabilistic forecasting approaches based on quantile regression methods Lago et al.
+(2021).
 However, these existing methods often focus on single-region markets, thereby neglecting the rich spatial information.
-In recent years, foundation models for time series forecasting have achieved remarkable success across diverse domains, demonstrating strong generalization capabilities by capturing complex data patterns (Zhou et al.
-2022; Wu et al.
-2023; Wang et al.
-2024), which makes them attractive candidates for electricity price forecasting.
-However, existing models are primarily designed for general univariate or multivariate time series tasks and are not explicitly tailored to electricity markets (Liang et al.
-2024).
-Moreover, some foundation models only provide pointwise forecasts and thus fail to capture the uncertainty essential for risk-aware trading decisions in volatile markets.
-Therefore, addressing the spatial interdependencies and uncertainty in electricity price forecasting within the integrated European market necessitates a tailored foundation model.
-To support the development of foundation models for electricity price forecasting, there is a pressing need for high-quality, large-scale, and up-to-date datasets that reflect the spatiotemporal complexity of the integrated European market.
-However, existing datasets are often fragmented in structure, cover only short time periods, are outdated, or focus on individual regions (Lago et al.
-2021).
-This lack of standardized, cross-border data poses a significant barrier to training and evaluating foundation models.
-In this paper, we bridge the gap between recent advances in foundation models and the requirements of electricity price forecasting in spatially interconnected markets.
-To this end, we introduce a comprehensive and up-to-date dataset capturing key spatiotemporal drivers of electricity prices across Europe.
-As a step forward, we propose PriceFM, a foundation model that incorporates graph-based inductive biases to learn spatial structures and generate probabilistic forecasts tailored to the integrated European electricity market.
+In recent years, foundation models for time series forecasting have achieved remarkable success across diverse domains, demonstrating strong generalization capabilities by capturing complex data patterns Zhou et al.
+(2022); Liu et al.
+(2023); Nie et al.
+(2023); Wu et al.
+(2023); Wang et al.
+(2024), which makes them attractive candidates for electricity price forecasting.
+In contrast to the conventional notion of foundation models based on pretraining, these models derive their foundation from a generic architecture applicable across domains, and are trained from scratch on the target data Liang et al.
+(2024).
+However, existing models are primarily designed for general univariate or multivariate time series tasks and are not explicitly tailored to electricity markets.
+Moreover, some foundation models only provide pointwise forecasts and thus fail to capture the uncertainty essential for risk-aware trading decisions.
+Therefore, addressing the spatial interdependencies and uncertainty in electricity price forecasting necessitates a tailored foundation model.
+To support the development of foundation models for electricity price forecasting, there is a pressing need for high-quality, large-scale, and up-to-date datasets that reflect the spatiotemporal complexity of integrated European markets.
+However, existing datasets are often fragmented in structure, cover only short time periods, are outdated, or focus on individual regions Lago et al.
+(2021).
+This lack of standardized data poses a significant barrier to training and evaluating foundation models.
+In this paper, we introduce a comprehensive and up-to-date dataset and propose PriceFM, a foundation model that incorporates graph-based inductive biases to generate probabilistic forecasts.
+Similar to other time-series foundation models, PriceFM adopts a generic architecture that can be trained from scratch and applied across various markets.
+Our contributions are as follows:
 Contribution
   • We introduce and release a comprehensive, up-to-date dataset.
-To the best of our knowledge, this is the largest and most diverse open dataset for European electricity markets, comprising day-ahead electricity prices, load forecasts, and solar and wind power generation forecasts (onshore and offshore), covering 24 European countries (38 regions), spanning from 2022-01-01 to 2025-01-01.
+To the best of our knowledge, this is the largest and most diverse open dataset for European electricity markets, comprising day-ahead electricity prices, day-ahead forecasts of load, solar, and wind power generation (onshore and offshore), covering 24 European countries (38 regions), spanning from 2022-01-01 to 2025-01-01.
   • We propose and release the PriceFM, a novel forecasting framework that integrates prior graph knowledge derived from the spatial topology of the European electricity market.
-PriceFM supports multi-region, multi-timestep, and multi-quantile probabilistic electricity price forecasting.
-  • We conduct extensive experiments to evaluate the model’s performance against multiple baselines, and assess the impact of design choices through ablation studies, thereby providing both quantitative evidence of overall performance and empirical insights into optimal configurations.
-Preliminary
+PriceFM supports joint multi-region, multi-timestep, and multi-quantile forecasting.
+  • We conduct experiments to evaluate the model’s performance against multiple baselines, and assess the impact of design choices through ablation studies, thereby providing both quantitative evidence of overall performance and insights into optimal configurations.
+2 Preliminary
 The forecasting target is $\mathcal{T}=24$ hourly prices for the delivery day $\mathcal{D}+1$, using data available before gate closure, typically around midday on day $\mathcal{D}$.
@@
 L=24{L}=24 corresponds to 24 hours from $\mathcal{D}$), for known electricity prices, denoted as $\mathbf{X}^{\mathrm{price}}_{r_{\mathrm{in}}}$.
-We also include forward-looking exogenous features, such as day‑ahead forecasts of load, solar, and wind (onshore and offshore) power generation for $\mathcal{D}+1$, denoted as $\mathbf{X}^{\mathrm{exo}}_{r_{\mathrm{in}}}$, made on $\mathcal{D}$ before gate closure, as well as their historical values over L{L}.
-The forecasting setup and the choice of feature set are widely used in prior works (Maciejowska 2020; Uniejewski and Weron 2021; Meng et al.
-2024), illustrated in Figure 2.
-Note that this work aims to utilize multi-region inputs to produce multi-region, multi-timestep, and multi-quantile probabilistic electricity price forecasts.
+We also include forward-looking exogenous features, such as day-ahead forecasts of load, solar, and wind (onshore and offshore) power generation for $\mathcal{D}+1$, denoted as $\mathbf{X}^{\mathrm{exo}}_{r_{\mathrm{in}}}$, made on $\mathcal{D}$ before gate closure, as well as their historical values over L{L}.
+The forecasting setup and the choice of feature set are widely used in prior works Maciejowska (2020); Uniejewski & Weron (2021); Meng et al.
+(2024).
+Importantly, this work aims to utilize multi-region inputs to produce multi-region, multi-timestep, and multi-quantile forecasts.
 Therefore, the input and output are defined as:
@@
 where $r_{\mathrm{in}},r_{\mathrm{out}}\in\mathcal{R}=\{\text{AT},\dots,\text{SK}\}$ (region codes, detailed in Appendix, Table 1.), $\tau\in\mathcal{Q}=\{0.1,0.5,0.9\}$ (quantile levels), $f_{1}=1$, and $f_{2}$ varies by region.
-For example, the forecasted wind power generation is not available in SI.
-Figure 2: Forecasting setup and feature set.
-Figure 3: European-level energy data in 2024, averaged across all available regions (depending on data availability).
-a Electricity price.
-Price spikes sharply during the morning and evening peak, dip around midday, and shows higher volatility in winter.
-b Forecasted load.
-Load exhibits a clear double‐peak each day (in morning and evening) with winter peaks substantially larger than summer.
-c Forecasted solar power generation.
+Figure 2: European-level energy data in 2024, averaged across regions.
+(a) Electricity price.
+Price spikes sharply during the morning and evening peak, dip around midday, and shows higher volatility in the second half of 2024.
+(b) Forecasted load.
+Load exhibits a double‐peak each day with winter peaks substantially larger than summer.
+(c) Forecasted solar power generation.
 Solar is zero overnight, rises in a smooth bell curve to a strong midday maximum, then falls back to zero by dusk, and is much higher in summer.
-d Forecasted wind power generation (sum of onshore and offshore).
-Wind lacks a daily pattern, fluctuates with high‐frequency spikes, and is much higher in the windy winter.
-Data
-The dataset is primarily collected from the European Network of Transmission System Operators for Electricity (ENTSO-E). We release the raw and cleaned dataset.
-Spatiotemporal Coverage
+(d) Forecasted wind power generation (onshore and offshore).
+Wind lacks a daily pattern, fluctuates with high‐frequency spikes, and is much higher in winter.
+3 Data
+3.1 Spatiotemporal Coverage
 Spatially, the dataset covers 24 European countries (38 regions).
@@
 For example, DK is split into two regions: DK1 and DK2.
-Each is connected to different neighboring regions, resulting in distinct cross-border power flows.
+Each is connected to different regions, resulting in distinct cross-border power flows.
 Temporally, the dataset spans from 2022-01-01 to 2025-01-01, providing wide temporal coverage.
-Feature Set
+3.2 Feature Set
 The feature set includes day-ahead electricity prices, load forecasts, and solar and wind power generation forecasts (onshore and offshore).
 For simplicity, we refer to these features as price, load, solar, and wind (onshore and offshore), respectively.
-Importantly, some regions only contain partial features.
-The full availability of features across regions is detailed in Appendix, Table 1.
+The availability of features across regions is detailed in Appendix, Table 1.
 A European-level visualization of these features is shown in Figure 2.
-Resolution
+3.3 Resolution
 We resample all features in an hourly resolution, as the raw data exhibit a heterogeneous temporal structure.
 For example, load from ES is provided at an hourly resolution before 2022-05-23 and then switches to a quarter-hourly resolution afterward; the price from AT is reported hourly, while the load is reported quarter-hourly.
-Missing Value
+3.4 Missing Value
 Partial features are excluded due to the high rate (above 15%) of missing values, summarized in Appendix, Table 1.
-For example, wind offshore from FR has a missing rate of 53.2% and is only available after 2023-08-07; solar from LV has a 75.5% missing rate and is only available after 2024-04-07; load from SK has a missing rate of 16.8% and is no longer available after 2024-07-01.
-For features with low missing rates (around 1%), such as load from EE and SI, we fill them using linear interpolation.
+For example, wind offshore from FR has a missing rate of 53.2% and is only available after 2023-08-07; load from SK has a missing rate of 16.8% and is no longer available after 2024-07-01.
+The features with low missing rates (below 1%) are filled using linear interpolation.
 Figure 3: Structure of PriceFM.
@@
 These regional representations are then stacked to form the spatial representation $\mathbf{S}$.
-Next, $\mathbf{S}$ is passed to the graph block to produce the representation $\mathbf{U}_{r_{\mathrm{out}}}$, which incorporates prior graph knowledge.
-Finally, these prior-knowledge-injected representations are fed into hierarchical quantile heads to produce multi-region, multi-timestep, and multi-quantile predictions.
-Model
-PriceFM, illustrated in Figure 4, aims to produce joint forecasts for multi-region, multi-timestep, and multi-quantile settings.
-The inputs $\mathbf{X}^{\mathrm{price}}_{r_{\mathrm{in}}}\in\mathbb{R}^{L\times f_{1}}$ and $\mathbf{X}^{\mathrm{exo}}_{r_{\mathrm{in}}}\in\mathbb{R}^{(L+\mathcal{T})\times f_{2}}$ for input region $r_{\mathrm{in}}$} are described in Section Preliminary.
-The inclusion of price and exogenous features introduces heterogeneity along the temporal and feature dimensions.
-Backbone
-We design a Fusion Block to enable effective fusion of heterogeneous multivariate input features and produce a spatial representation, and a Graph Block to incorporate prior graph knowledge to enforce spatial decay.
-Fusion Block
-We first align the temporal dimensions by projecting the price feature to length $L+\mathcal{T}$ via a dense layer with linear activation, followed by a transformation into a hidden space of dimension $h$ via $k$ dense layers:
+Next, $\mathbf{S}$ is passed to the graph block to produce the spatial representation $\mathbf{U}_{r_{\mathrm{out}}}$.
+Finally, $\mathbf{U}_{r_{\mathrm{out}}}$ is fed into hierarchical quantile heads to produce joint forecasts.
+4 Model
+PriceFM, illustrated in Figure 3, aims to produce multi-region, multi-timestep, and multi-quantile forecasts.
+The inputs $\mathbf{X}^{\mathrm{price}}_{r_{\mathrm{in}}}\in\mathbb{R}^{L\times f_{1}}$ and $\mathbf{X}^{\mathrm{exo}}_{r_{\mathrm{in}}}\in\mathbb{R}^{(L+\mathcal{T})\times f_{2}}$ are described in Section Preliminary, introducing heterogeneity along the temporal and feature dimensions.
+4.1 Fusion Block
+We first project the temporal dimensions by transposing and projecting $\mathbf{X}^{\mathrm{price}}_{r_{\mathrm{in}}}$ to length $L+\mathcal{T}$ via a dense layer with linear activation, and then transpose back and project the feature dimension of both price and exogenous features into a hidden space of dimension $h$ via $k$ dense layers:
 $\mathbf{X}^{\mathrm{price}}_{r_{\mathrm{in}}}$ $\xrightarrow{\mathrm{Project}}\tilde{\mathbf{X}}^{\mathrm{price}}_{r_{\mathrm{in}}}\in\mathbb{R}^{(L+\mathcal{T})\times f_{1}},$ (1)
-$\tilde{\mathbf{X}}^{\mathrm{price}}_{r_{\mathrm{in}}}$ $\xrightarrow{\mathrm{Dense}\ast k}\hat{\mathbf{X}}^{\mathrm{price}}_{r_{\mathrm{in}}}\in\mathbb{R}^{(L+\mathcal{T})\times h},$ (2)
-${\mathbf{X}}^{\mathrm{exo}}_{r_{\mathrm{in}}}$ $\xrightarrow{\mathrm{Dense}\ast k}\hat{\mathbf{X}}^{\mathrm{exo}}_{r_{\mathrm{in}}}\in\mathbb{R}^{(L+\mathcal{T})\times h}.$ (3)
+$\tilde{\mathbf{X}}^{\mathrm{price}}_{r_{\mathrm{in}}}$ $\xrightarrow{\mathrm{Project}}\hat{\mathbf{X}}^{\mathrm{price}}_{r_{\mathrm{in}}}\in\mathbb{R}^{(L+\mathcal{T})\times h},$ (2)
+${\mathbf{X}}^{\mathrm{exo}}_{r_{\mathrm{in}}}$ $\xrightarrow{\mathrm{Project}}\hat{\mathbf{X}}^{\mathrm{exo}}_{r_{\mathrm{in}}}\in\mathbb{R}^{(L+\mathcal{T})\times h}.$ (3)
 Next, we perform feature fusion through residual addition and flatten the fused representation to produce the regional representation $\mathbf{F}_{r_{\mathrm{in}}}$, which encodes the backward-looking price feature and forward-looking contextual information:
@@
 $\mathbf{S}=\mathrm{Stack}\left(\{\mathbf{F}_{r_{\mathrm{in}}}\}_{{r_{\mathrm{in}}}\in\mathcal{R}}\right)\in\mathbb{R}^{|\mathcal{R}|\times(L+\mathcal{T})\cdot h}.$ (5)
-This fusion process enables feature aggregation across input regions with varying temporal and feature dimensions.
-Graph Block
-We first construct graph distance by performing a breadth-first search (BFS) traversal on the cross-border grid topology, which records the prior graph knowledge.
-For a given output region $r_{\mathrm{out}}\in\mathcal{R}$, we define the graph distance $d(r_{\mathrm{in}},r_{\mathrm{out}})$ as the minimal number of transmission hops from each input region $r_{\mathrm{in}}$ to out region $r_{\mathrm{out}}$, based on direct or indirect physical connectivity:
+4.2 Graph Block
+We first construct graph distance by performing a breadth-first search (BFS) traversal on the cross-border grid topology, detailed in Appendix, Table 6.
+For a given output region $r_{\mathrm{out}}\in\mathcal{R}$, we define the graph distance $d(r_{\mathrm{in}},r_{\mathrm{out}})$ as the minimal number of transmission hops from each input region $r_{\mathrm{in}}$ to the output region $r_{\mathrm{out}}$, based on direct or indirect physical connectivity:
 $d(r_{\mathrm{in}},r_{\mathrm{out}})=\begin{cases}0&\text{if }r_{\mathrm{in}}=r_{\mathrm{out}},\\ 1&\text{if }r_{\mathrm{in}}\sim r_{\mathrm{out}},\\ 1+\min\limits_{r^{\prime}\sim r_{\mathrm{in}}}d(r^{\prime},r_{\mathrm{out}})&\text{otherwise},\end{cases}$ (6)
@@
 The region HU is directly connected to AT, thus $d(\text{HU},\text{AT})=1$.
-SK is not directly connected to AT but directly connected to HU, yielding $d(\text{SK},\text{AT})=2$.
-Next, we introduce a graph decay mechanism to incorporate spatial prior knowledge.
+SK is indirectly connected to AT via HU, yielding $d(\text{SK},\text{AT})=2$.
+Next, we introduce a graph decay mechanism to inject prior graph knowledge.
 Intuitively, input regions that are topologically closer to the output region $r_{\mathrm{out}}$ are expected to exert a stronger influence than more distant ones.
@@
 For $c<0$ (Regime III), the decay weight decreases more gradually, preserving the influence of neighbors.
-Using this decay function, the decay mask is constructed as:
+We construct the decay mask as:
 $\mathbf{W}_{r_{\mathrm{out}}}=\begin{bmatrix}w(d(\text{AT},r_{\mathrm{out}});c;D)\\ w(d(\text{BE},r_{\mathrm{out}});c;D)\\ \vdots\\ w(d(\text{SK},r_{\mathrm{out}});c;D)\end{bmatrix}\in\mathbb{R}^{|\mathcal{R}|\times 1}.$ (8)
@@
 $\mathbf{U}_{r_{\mathrm{out}}}=\frac{\mathbf{W}_{r_{\mathrm{out}}}^{\top}\mathbf{S}}{\mathbf{W}_{r_{\mathrm{out}}}^{\top}\mathbf{1}},$ (9)
-where $\mathbf{W}_{r_{\mathrm{out}}}\in\mathbb{R}^{|\mathcal{R}|\times 1}$ is the decay mask vector for $r_{\mathrm{out}}$, and $\mathbf{1}\in\mathbb{R}^{|\mathcal{R}|\times 1}$ is a vector of ones.
-This operation acts as spatial regularization and eliminates the need for an exhaustive learning process to determine spatial weights, as required in methods such as attention.
-Head
+where $\mathbf{1}\in\mathbb{R}^{|\mathcal{R}|\times 1}$ is a vector of ones.
+This operation acts as spatial regularization and eliminates the need for an exhaustive learning process to determine spatial weights, as required in methods such as the attention mechanism.
+4.3 Head
 We design a multi-region, multi-timestep, and multi-quantile head, where the model produces joint probabilistic forecasts.
 To prevent quantile crossing issue [^1] Quantile crossing refers to the phenomenon where upper quantile predictions (e.g., 90%) fall below lower quantiles (e.g., 10%), violating the monotonicity of the quantile function.
-(Chernozhukov, Fernández-Val, and Galichon 2010)., we adopt a hierarchical quantile head (Yu et al.
-2025).
+Chernozhukov et al.
+(2010)., we adopt a hierarchical quantile head Yu et al.
+(2025).
 In detail, the median quantile ($\tau=0.5$) price trajectory, which represents the full set of timesteps $\mathcal{T}$, is predicted from $\mathbf{U}_{r_{\mathrm{out}}}$ via a dense layer:
 $\hat{\mathbf{y}}_{{r_{\mathrm{out}}},0.5}=\text{Dense}(\mathbf{U}_{r_{\mathrm{out}}})\in\mathbb{R}^{\mathcal{T}}.$ (10)
-For the upper quantile ($\tau=0.9$), a residual price trajectory $\hat{\mathbf{r}}_{{r_{\mathrm{out}}},0.9}$ is predicted from $\mathbf{U}_{r_{\mathrm{out}}}$ via a dense layer:
+For the upper quantile ($\tau=0.9$), a residual price trajectory $\hat{\mathbf{r}}_{{r_{\mathrm{out}}},0.9}$ is produced from $\mathbf{U}_{r_{\mathrm{out}}}$:
 $\hat{\mathbf{r}}_{{r_{\mathrm{out}}},0.9}=\text{Dense}(\mathbf{U}_{r_{\mathrm{out}}})\in\mathbb{R}^{\mathcal{T}},$ (11)
-where a non-negative function $g(\cdot)$, such as absolute-value function, is applied to the price residual.
-The upper quantile forecast is obtained by adding this non-negative residual to the median prediction:
+where a non-negative function $g(\cdot)$, such as absolute-value function, is applied to the price residual. The upper quantile forecast is obtained by adding this non-negative residual to the median:
 $\hat{\mathbf{y}}_{{r_{\mathrm{out}}},0.9}=\hat{\mathbf{y}}_{{r_{\mathrm{out}}},0.5}+g(\hat{\mathbf{r}}_{{r_{\mathrm{out}}},0.9}).$ (12)
@@
 $\hat{\mathbf{y}}_{{r_{\mathrm{out}}},0.1}=\hat{\mathbf{y}}_{{r_{\mathrm{out}}},0.5}-g(\hat{\mathbf{r}}_{{r_{\mathrm{out}}},0.1}).$ (14)
-This hierarchical design guarantees that the upper quantile prediction is greater than or equal to the lower one at each time step, effectively overcoming quantile crossing.
-Loss
+This hierarchical design guarantees that the upper quantile prediction is greater than or equal to the lower one at each time step, overcoming quantile crossing.
+4.4 Loss
 We introduce the Average Quantile Loss (AQL) as the training objective for multi-region, multi-timestep, and multi-quantile probabilistic forecasting.
 Let $y_{i,{r_{\mathrm{out}}},t}$ denote the ground-truth price for the $i$-th training sample, output region $r_{\mathrm{out}}$}, and timestep $t$, and let $\hat{y}_{i,{r_{\mathrm{out}}},t,\tau}$ be the corresponding predicted quantile.
-The AQL across all training samples, regions, timesteps, and quantiles is computed as:
+The AQL is computed as:
 $\mathrm{AQL}=\frac{1}{N\left|\mathcal{R}\right|\mathcal{T}\left|\mathcal{Q}\right|}\sum_{i=1}^{N}\sum_{r_{\mathrm{out}}\in\mathcal{R}}\sum_{t=1}^{\mathcal{T}}\sum_{\tau\in\mathcal{Q}}L_{\tau}\left(y_{i,r_{\mathrm{out}},t},\hat{y}_{i,r_{\mathrm{out}},t,\tau}\right),$ (15)
-where $N$ is the number of samples, and the quantile loss function $L_{\tau}$ for a single prediction task is defined as:
+where $N$ is the number of samples, and the quantile loss $L_{\tau}$ is defined as:
 $L_{\tau}(y,\hat{y}_{\tau})=\begin{cases}\tau\cdot(y-\hat{y}_{\tau}),&\text{if }y\geq\hat{y}_{\tau},\\ (1-\tau)\cdot(\hat{y}_{\tau}-y),&\text{otherwise},\end{cases}$ (16)
 where $y$ and $\hat{y}$ are the true and predicted values, respectively.
-A lower AQL indicates more accurate probabilistic forecasting across spatial and temporal dimensions.
-Other Details
-We use the Adam optimizer with an initial learning rate of $4\times 10^{-3}$, which decays exponentially by a factor of 0.95 every 10 epochs.
-The batch size is 8, which introduces a slight regularization effect.
-The model is trained for 50 epochs, and the checkpoint with the lowest validation loss is saved.
-Experiment
-For the three-year dataset, we use the first two years of the dataset (67%) for training, and the final year (33%) is evenly split into validation and testing sets, strictly preserving the temporal order to ensure realistic out-of-sample evaluation.
-We assess the model performance using the quantile loss at individual levels (Q10, Q50, and Q90), AQL, Average Quantile Crossing Rate (AQCR), Root Mean Squared Error (RMSE), Mean Absolute Error (MAE), and coefficient of determination ($R^{2}$).
-The Diebold-Mariano (DM) test is applied to determine if two models have a significant difference in the probabilistic and pointwise forecasting (Diebold and Mariano 2002).
-Each metric is explained in Appendix, Section Metrics.
-Model Comparison
-We include several spatial and temporal models as baselines. Model descriptions and hyperparameter optimization are detailed in Appendix, Section Baselines.
-Spatial Models
-Following (Yang et al.
-2024; Meng et al.
-2024), where variants of GNNs, such as the Graph Convolutional Network (GraphConv) and the Graph Attention Network (GraphAttn), have made successes in electricity price forecasting, we include them as baselines.
-We also compare with other advanced variants, such as a GNN based on sample-and-aggregate operations (GraphSAGE) (Hamilton, Ying, and Leskovec 2017), a diffusion-based GNN (GraphDiff) (Li et al.
-2018), and a GNN with Auto-Regressive Moving Average filters (GraphARMA) (Bianchi et al.
-2021).
-Results in Table 1 show that PriceFM statistically outperforms all spatial models, confirmed by both the probabilistic and pointwise DM test.
-A common limitation of these models is that they have only limited spatial priors, relying primarily on data-driven learning to assign spatial importance.
-Without an explicit mechanism to enforce spatial decay, they lack regularization over noisy regions.
+5 Experiment
+We split the data into training (2022-01-01 to 2024-01-01), validation (2024-01-01 to 2024-07-01), and testing (2024-07-01 to 2025-01-01).
+The choice of the testing period aims to include numerous extreme prices, as illustrated in Figure 2 (a).
+We assess the model performance using the quantile losses ($\text{Q}_{0.1}$, $\text{Q}_{0.5}$, and $\text{Q}_{0.9}$), AQL, Average Quantile Crossing Rate (AQCR), Root Mean Squared Error (RMSE), Mean Absolute Error (MAE), and Coefficient of Determination ($\text{R}^{2}$).
+The Diebold-Mariano (DM) test is applied to determine if two models have a significant difference Diebold & Mariano (2002).
+All metrics are explained in Appendix G.
+The hyperparameters are detailed in Appendix D
+5.1 Model Comparison
+5.1.1 Naïve Baselines
+We include three seasonal naïve baselines as reference models, where only historical prices are used as input: Naïve1 uses 24 prices from the previous day; Naïve2 uses 24 prices averaged over the past three days; Naïve3 uses 24 prices averaged over the past seven days.
+To obtain probabilistic results, we compute empirical quantiles at individual levels ($\mathcal{Q}=\{0.1,0.5,0.9\}$) for each delivery hour.
+The seasonal naïves are commonly used to evaluate the autoregressive strength of the signal and often serve as strong baselines Ziel & Weron (2018); Lago et al.
+(2021).
+The results from Table 1 show that PriceFM significantly outperforms the naïve baselines, confirmed by both the probabilistic and pointwise DM tests, with all $p\text{-values}<0.05$ and negative DM values.
+Specifically, the AQL values of baselines are between 36.83% and 44.80% higher, with an AQCR of 0.00%, as these forecasts are directly computed from historical values.
+Moreover, the high RMSE and MAE, together with the low $\text{R}^{2}$, observed in the naïve baselines suggest limited performance.
+5.1.2 Graph Models
+We compare with multiple GNN variants: Graph Convolutional Network (GCN) Kipf (2016), Graph Attention Network (GAT) Veličković et al.
+(2017), GraphSAGE Hamilton et al.
+(2017), GraphDiffusion Li et al.
+(2018), and GraphARMA Bianchi et al.
+(2021).
+The adjacency matrix of these graph models is explained in Appendix E.
+From Table 1, we observe that PriceFM outperforms all graph models, confirmed by both the probabilistic and pointwise DM tests, with all $p\text{-values}<0.05$ and negative DM values.
+Notably, PriceFM reduces the AQL by between 8.86% and 15.77% and consistently achieves 0.00% AQCR compared to graph baselines.
+A common limitation of these graph baselines is that they lack explicit regularization over noisy regions and rely primarily on data-driven learning to assign spatial importance.
 As a result, they require large datasets to generalize.
-However, despite our dataset spanning three years, which is considered large in the domain, the daily forecasting requirement limits the training set to only around 700 samples, making it unsuitable for such models.
-Model Q10 $\downarrow$ Q50 $\downarrow$ Q90 $\downarrow$ AQL $\downarrow$ AQCR $\downarrow$ RMSE $\downarrow$ MAE $\downarrow$ ${R}^{2}$ $\uparrow$
-GraphConv 4.81±0.16 10.76±0.16 7.11±0.33 7.56±0.15 2.19±0.76 35.75±0.60 21.53±0.33 0.51±0.02
-GraphAttn 5.08±0.31 11.58±0.37 7.90±0.40 8.18±0.29 1.41±0.47 37.63±0.62 23.15±0.75 0.44±0.01
-GraphSAGE 5.20±0.13 11.27±0.23 7.39±0.35 7.95±0.18 2.79±0.56 37.09±0.81 22.53±0.46 0.47±0.02
-GraphDiff 4.80±0.20 11.03±0.21 7.34±0.33 7.73±0.18 3.33±0.70 36.35±0.80 22.07±0.42 0.48±0.02
-GraphARMA 4.87±0.07 11.10±0.22 7.00±0.22 7.66±0.16 2.05±0.65 36.15±0.59 22.21±0.44 0.49±0.02
-FEDFormer 47.55±1.51 28.91±0.99 0.14±0.00
-TimesNet 44.20±0.87 26.40±0.52 0.30±0.01
-TimeXer 47.38±1.18 28.29±0.71 0.14±0.02
-PriceFM 4.89±0.05 9.83±0.15 5.96±0.08 6.89±0.08 0.00±0.00 32.25±0.38 19.65±0.30 0.61±0.01
-Table 1: Model performance comparison on the testing set.
+Despite our dataset spanning three years, which is considered large in the domain, the daily forecasting requirement limits the training set to only around 700 samples, making it unsuitable for such models.
+Table 1: Model comparison on the testing set.
 The symbol “–” indicates that the model does not support probabilistic forecasting by design.
+The naïve baselines have zero standard deviation, as they are deterministically computed.
 All metrics are reported as mean±standard deviation over 5 independent runs.
 The best results are shown in bold, and the second-best are underlined.
-The units of Q10, Q50, Q90, AQL, RMSE, and MAE are expressed in €/MWh (denormalized), and AQCR in %.
-Temporal Models
-We include FEDFormer (Zhou et al.
-2022), TimesNet (Wu et al.
-2023), and TimeXer (Wang et al.
-2024) in our evaluation.
+The units of $\text{Q}_{0.1}$, $\text{Q}_{0.5}$, $\text{Q}_{0.9}$, AQL, RMSE, and MAE are expressed in €/MWh, and AQCR in %.
+Model $\text{Q}_{\textbf{{0.1}}}$ $\downarrow$ $\text{Q}_{\textbf{{0.5}}}$ $\downarrow$ $\text{Q}_{\textbf{{0.9}}}$ $\downarrow$ AQL $\downarrow$ AQCR $\downarrow$ RMSE $\downarrow$ MAE $\downarrow$ $\textbf{R}^{\textbf{2}}$ $\uparrow$
+Naïve1 5.68±0.00 14.30±0.00 8.31±0.00 9.43±0.00 0.00±0.00 46.90±0.00 28.60±0.00 0.16±0.00
+Naïve2 6.01±0.00 14.42±0.00 8.44±0.00 9.62±0.00 0.00±0.00 46.73±0.00 28.84±0.00 0.17±0.00
+Naïve3 6.65±0.00 14.58±0.00 8.71±0.00 9.98±0.00 0.00±0.00 46.73±0.00 29.15±0.00 0.17±0.00
+GCN 4.81±0.16 10.76±0.16 7.11±0.33 7.56±0.15 2.19±0.76 35.75±0.60 21.53±0.33 0.51±0.02
+GAT 5.08±0.31 11.58±0.37 7.90±0.40 8.18±0.29 1.41±0.47 37.63±0.62 23.15±0.75 0.44±0.01
+GraphSAGE 5.20±0.13 11.27±0.23 7.39±0.35 7.95±0.18 2.79±0.56 37.09±0.81 22.53±0.46 0.47±0.02
+GraphDiffusion 4.80±0.20 11.03±0.21 7.34±0.33 7.73±0.18 3.33±0.70 36.35±0.80 22.07±0.42 0.48±0.02
+GraphARMA 4.87±0.07 11.10±0.22 7.00±0.22 7.66±0.16 2.05±0.65 36.15±0.59 22.21±0.44 0.49±0.02
+FEDFormer 44.60±0.88 27.53±0.81 0.30±0.01
+PatchTST 45.32±1.03 26.21±0.92 0.29±0.02
+iTransformer 45.14±0.96 27.05±0.64 0.29±0.02
+TimesNet 44.20±0.87 26.40±0.52 0.30±0.01
+TimeXer 44.57±0.66 26.52±0.55 0.30±0.01
+PriceFM 4.80±0.06 9.81±0.17 5.96±0.08 6.89±0.12 0.00±0.00 32.24±0.39 19.68±0.31 0.61±0.01
+5.1.3 Time-Series Foundation Models
+We include several time-series foundation models: FEDFormer Zhou et al.
+(2022), iTransformer Liu et al.
+(2023), PatchTST Nie et al.
+(2023), TimesNet Wu et al.
+(2023), and TimeXer Wang et al.
+(2024), to investigate whether these pure time-series models can capture spatial patterns without prior graph knowledge.
 As these foundation models do not support graph-based input, features from all regions are concatenated along the feature dimension.
-Including these models aims to investigate whether these pure time-series models can capture spatial patterns without prior graph knowledge.
-Results in Table 1 show that PriceFM statistically outperforms all temporal models, confirmed by the pointwise DM test.
-In particular, the inclusion of noisy features from weakly related regions easily leads to overfitting, given the high model complexity of these time-series foundation models.
-This observation confirms that the temporal models struggle to recognize spatial patterns without explicit injection of graph knowledge.
+The results in Table 1 show that these time-series foundation models achieve similar RMSE and MAE but exhibit better explained variance, as indicated by higher $\text{R}^{2}$, compared to the naïve baselines.
+Notably, PriceFM outperforms all time-series foundation models, confirmed by the pointwise DM test, with all $p\text{-values}<0.05$ and negative DM values.
+On average, PriceFM improves RMSE, MAE, and $\text{R}^{2}$ by 27.98%, 26.38%, and 0.31, respectively.
+Given the high complexity of these foundation models, the inclusion of features from noisy regions easily leads to overfitting.
+This observation confirms that the pure time-series models struggle to recognize useful spatial patterns.
 In contrast, PriceFM incorporates a graph decay mechanism that acts as a spatial regularizer, thereby attenuating the influence of noisy regions.
-Ablation Study
-We systematically evaluate the impact of key hyperparameters and architectural design choices, including spatiotemporal configurations, fusion strategies in the fusion block, graph decay mechanisms in the graph block, and variants of the hierarchical quantile head.
-Spatiotemporal Configurations
-We assess the impact of two key hyperparameters that control the spatial decay strength and temporal input length:
+5.2 Ablation Study
+5.2.1 Spatiotemporal Configurations
   • Curvature Parameter: Spatially, we evaluate $c\in\{-1.0,-0.8,\ldots,0.8,1.0\}$ in increments of 0.2, ranging from weak decay to strong decay.
-In total, 2,090 trials are conducted to determine the optimal curvature value for each output region individually, based on validation loss.
+In total, 2,090 trials are conducted to determine the optimal curvature value for each output region individually.
   • Backward-Looking Window Size: Temporally, we compare $L\in\{24,72,168\}$, corresponding to one day, three days, and one week. For each window size, all other hyperparameters are re-optimized.
-Figure 5 illustrates the testing loss and the distribution of optimal curvature values across all regions.
+Spatially, Figure 4 illustrates the testing loss and the distribution of optimal curvature values across all regions.
 Most regions confirm spatial interdependencies ($c\neq 1.0$).
-The identified optimal curvature shows a weak negative correlation ($\rho=-0.18$) with the number of directly connected regions, mentioned in Figure 1.
-The results in Table 2 indicate that the optimal backward-looking window size is 24, potentially because information from the distant past becomes outdated.
+Temporally, the results in Table 2 indicate that the optimal backward-looking window size is 24, potentially because information from the distant past becomes outdated.
 Figure 4: Spatial distribution of testing loss and curvature values.
-a Average quantile loss per region on the testing set.
+(a) Average quantile loss per region on the testing set.
 Western European regions exhibit lower losses, whereas BG, HU, and RO show particularly high losses (orange areas).
-b Optimal curvature value per region.
-Particularly, 8 regions from IT and SE have a curvature value of 1, indicating optimal performance by excluding neighboring features.
-Method Q10 $\downarrow$ Q50 $\downarrow$ Q90 $\downarrow$ AQL $\downarrow$ AQCR $\downarrow$ RMSE $\downarrow$ MAE $\downarrow$ ${R}^{2}$ $\uparrow$
-$L=24^{\dagger}$ 4.89±0.05 9.83±0.15 5.96±0.08 6.89±0.08 0.00±0.00 32.25±0.38 19.65±0.30 0.61±0.01
+(b) Optimal curvature value per region.
+Notably, BG, ES, FR, FI, and partial regions from IT and SE have a curvature value of $1.0$, indicating optimal performance by excluding neighboring features.
+Table 2: Ablation studies of different module choices: temporal configuration (rows 1-3), fusion block (rows 4–6), graph block (rows 7–9), and hierarchical quantile head (rows 10–13).
+The symbol marks the method used in PriceFM.
+Method $\text{Q}_{\textbf{{0.1}}}$ $\downarrow$ $\text{Q}_{\textbf{{0.5}}}$ $\downarrow$ $\text{Q}_{\textbf{{0.9}}}$ $\downarrow$ AQL $\downarrow$ AQCR $\downarrow$ RMSE $\downarrow$ MAE $\downarrow$ $\textbf{R}^{\textbf{2}}$ $\uparrow$
+$L=24^{\dagger}$ 4.80±0.06 9.81±0.17 5.96±0.08 6.89±0.12 0.00±0.00 32.24±0.39 19.68±0.31 0.61±0.01
 $L=72$ 5.09±0.13 10.04±0.20 6.27±0.06 7.14±0.09 0.00±0.00 32.76±0.41 20.09±0.39 0.60±0.01
 $L=168$ 5.62±0.30 10.52±0.26 6.65±0.16 7.60±0.21 0.00±0.00 33.82±0.59 21.04±0.53 0.57±0.01
-Res. Add 4.89±0.05 9.83±0.15 5.96±0.08 6.89±0.08 0.00±0.00 32.25±0.38 19.65±0.30 0.61±0.01
+Res. Add 4.80±0.06 9.81±0.17 5.96±0.08 6.89±0.12 0.00±0.00 32.24±0.39 19.68±0.31 0.61±0.01
 Concat. 5.03±0.18 10.50±0.21 6.40±0.06 7.31±0.13 0.00±0.00 34.68±0.49 21.01±0.42 0.56±0.01
 Cross-Attn 4.92±0.14 9.76±0.10 6.07±0.09 6.92±0.09 0.00±0.00 32.46±0.27 19.51±0.19 0.61±0.01
-Decay 4.89±0.05 9.83±0.15 5.96±0.08 6.89±0.08 0.00±0.00 32.25±0.38 19.65±0.30 0.61±0.01
+Decay 4.80±0.06 9.81±0.17 5.96±0.08 6.89±0.12 0.00±0.00 32.24±0.39 19.68±0.31 0.61±0.01
 Random 5.32±0.27 10.95±0.13 7.21±0.11 7.83±0.11 0.00±0.00 35.85±0.30 21.89±0.25 0.50±0.02
 No Decay 5.23±0.32 11.05±0.22 6.84±0.22 7.71±0.16 0.00±0.00 35.81±0.63 22.11±0.44 0.50±0.02
-ABS 4.89±0.05 9.83±0.15 5.96±0.08 6.89±0.08 0.00±0.00 32.25±0.38 19.65±0.30 0.61±0.01
+ABS 4.80±0.06 9.81±0.17 5.96±0.08 6.89±0.12 0.00±0.00 32.24±0.39 19.68±0.31 0.61±0.01
 Square 4.86±0.14 9.80±0.16 6.05±0.21 6.90±0.15 0.00±0.00 32.35±0.36 19.59±0.33 0.61±0.01
@@
 Standard 4.80±0.14 10.05±0.06 5.94±0.13 6.93±0.04 4.10±1.28 32.76±0.18 20.10±0.13 0.60±0.01
-Table 2: Ablation studies of different module choices: temporal configuration (rows 1-3), fusion block (rows 4–6), graph block (rows 7–9), and hierarchical quantile head (rows 10–13).
-The symbol marks the method used in PriceFM.
-All metrics are reported as mean±standard deviation over 5 independent runs.
-The best results are shown in bold, and the second-best are underlined.
-The units of Q10, Q50, Q90, AQL, RMSE, and MAE are expressed in €/MWh (denormalized), and AQCR in %.
-Fusion Block
-We compare the residual addition from the fusion block, described in Eq. (4), with two fusion strategies:
-  • Concatenation: We replace Eq. (4) by first flattening both $\hat{\mathbf{X}}^{\mathrm{price}}_{r_{\mathrm{in}}}$ and $\hat{\mathbf{X}}^{\mathrm{exo}}_{r_{\mathrm{in}}}$, then concatenating them:
+5.2.2 Fusion Block
+  • Concatenation: We replace Equation 4 by first flattening both $\hat{\mathbf{X}}^{\mathrm{price}}_{r_{\mathrm{in}}}$ and $\hat{\mathbf{X}}^{\mathrm{exo}}_{r_{\mathrm{in}}}$, and then concatenating them:
     $\mathbf{F}_{r_{\mathrm{in}}}=\mathrm{Concat}\left(\mathrm{Flatten}\left(\hat{\mathbf{X}}^{\mathrm{price}}_{r_{\mathrm{in}}}\right),\,\mathrm{Flatten}\left(\hat{\mathbf{X}}^{\mathrm{exo}}_{r_{\mathrm{in}}}\right)\right).$ (17)
@@
 The results in Table 2 show that replacing the residual addition with concatenation leads to 6.10% higher AQL for probabilistic prediction.
-Switching to cross-attention yields comparable performance to residual addition, while introducing additional model parameters and unnecessary complexity.
-This suggests that the residual addition strikes a favorable balance between performance and model simplicity.
-Graph Block
-We compare the graph decay mask from the graph block, described in Eq. (8), with two alternatives:
-  • Random Graph Decay Mask: We replace Eq.
-(8) with a randomly sampled vector, where each decay weight is drawn independently from a uniform distribution over $[0,1]$, thereby removing the spatial graph prior:
+Switching to cross-attention yields comparable performance to residual addition, while introducing additional model parameters.
+This suggests that the residual addition strikes a favorable balance between predictive performance and model simplicity.
+5.2.3 Graph Block
+  • Random Graph Decay Mask: We replace Equation 8 with a randomly sampled vector, where each decay weight is drawn independently from a uniform distribution over $[0,1]$, thereby removing the spatial graph prior:
     $\mathbf{W}_{r_{\mathrm{out}}}\sim\mathcal{U}(0,1)^{|\mathcal{R}|\times 1}.$ (19)
-  • No Graph Decay: We remove the decay mask, which simplifies Eq. (9) to a uniform average over input regions:
+  • No Graph Decay: We remove the decay mask, which simplifies Equation 9 to a uniform average over input regions:
     $\mathbf{U}_{r_{\mathrm{out}}}=\frac{\mathbf{1}^{\top}\mathbf{S}}{|\mathcal{R}|},$ (20)
 The results in Table 2 demonstrate that randomizing or removing the graph decay mask, which discards the prior graph knowledge, leads to a significant drop in all metrics.
-We also observe that such results, obtained without prior graph knowledge, are on par with those of GNN variants, shown in Table 1.
+We also observe that such results are on par with those of GNN baselines.
 We emphasize that relying on pure data-driven learning without an explicit decay mechanism leads to a loss of the key inductive bias, limiting the model’s performance, especially when the training data is scarce.
-Hierarchical Quantile Head
-We evaluate the impact of different non-negative functions used in Eq. (12) and (14), and include a standard multi-quantile head for comparison.
-  • Non-Negative Functions: We replace the absolute-value function with either a square function or ReLU:
+5.2.4 Hierarchical Quantile Head
+  • Non-Negative Functions: We replace the absolute-value function used in Equation 12 and 14 with either a square function or ReLU:
     $g(\cdot)=(\cdot)^{2},$ (21)
     $g(\cdot)=\max(0,\cdot).$ (22)
-  • Standard Multi-Quantile Head: The Eq.
-(11) and (13) are skipped, and the shared representation $\mathbf{U}_{r_{\mathrm{out}}}$ is passed directly to three independent dense layers to produce each quantile trajectory:
-    $\hat{\mathbf{y}}_{r_{\mathrm{out}},0.1}=\mathrm{Dense}(\mathbf{U}_{r_{\mathrm{out}}})\in\mathbb{R}^{\mathcal{T}},$ (23)
-    $\hat{\mathbf{y}}_{r_{\mathrm{out}},0.5}=\mathrm{Dense}(\mathbf{U}_{r_{\mathrm{out}}})\in\mathbb{R}^{\mathcal{T}},$ (24)
-    $\hat{\mathbf{y}}_{r_{\mathrm{out}},0.9}=\mathrm{Dense}(\mathbf{U}_{r_{\mathrm{out}}})\in\mathbb{R}^{\mathcal{T}}.$ (25)
+  • Standard Multi-Quantile Head: The Equation 11 and 13 are skipped, and $\mathbf{U}_{r_{\mathrm{out}}}$ is passed directly to three independent dense layers to produce quantile trajectories.
 The results in Table 2 reveal that replacing the absolute-value function with either a square function or ReLU does not result in a noticeable change in performance, suggesting that the choice of non-negative function is flexible.
 Moreover, while the hierarchical quantile head achieves comparable loss to the standard multi-quantile head, the latter exhibits a mean AQCR of 4.10%, indicating that the hierarchical design mitigates quantile crossing without harming performance.
-Conclusion
+6 Conclusion
 In this paper, we introduced and released a dataset, which will benefit both the research community and the energy industry.
-We proposed and released PriceFM, a foundation model tailored for interconnected European electricity markets.
+We proposed and released PriceFM, a foundation model with a generic architecture applicable to all European electricity markets without relying on pretraining, similar to time-series foundation models used in this work.
 Extensive experiments and ablation studies demonstrate that PriceFM outperforms competitive baselines and highlight the importance of spatial context.
-By enabling more accurate and comprehensive probabilistic electricity price forecasting, our work has the potential to support better decision-making in energy trading and grid management, contributing to a more efficient and resilient energy ecosystem.
-Despite these promising results, our study has two key limitations.
-First, the graph decay function is empirically designed; the performance may improve with alternative functions.
-Second, PriceFM relies on the current transmission topology, which may become outdated as transmission networks evolve, requiring retraining to adapt to new scenarios.
-Appendix
-Code Guideline
+By enabling more accurate and comprehensive probabilistic electricity price forecasting, our work has the potential to support better decision-making in energy trading and grid management.
+Several directions remain for future work.
+First, as the current graph decay function is empirically defined, exploring alternative formulations could improve spatial representations.
+Second, as the transmission network evolves, model retraining may be required to account for structural changes.
+Third, since our dataset is sampled at an hourly resolution, we aim to extend PriceFM to support quarter-hourly forecasts, in anticipation of more European bidding zones transitioning to 15-minute markets in the coming years.
+7 Reproducibility Statement
+To support reproducibility, we release well-documented code with an easy-to-use three-step pipeline, along with the code structure and usage guidelines in Appendix A.
+We report hardware and runtime details in Appendix B to facilitate realistic deployment of our proposed model.
+As the release and cleaning of the dataset are part of our contributions, the details are described in Section Data and Appendix C.
+Additional reproducibility details are also provided, including data scaling (Appendix F), evaluation metrics (Appendix G), hyperparameters and training procedures (Appendix D).
+Appendix A Code Guideline
 We open-source all code for preprocessing, modeling, and analysis. The project directory is structured as follows:
@@
    |- Data/
+   |- Figure/
    |- Model/
@@
    |- Main.py
+   |- Tutorial.ipynb
    |- README.md
@@
 To facilitate reproducibility and accessibility, we have streamlined the entire pipeline through extensive engineering efforts into just three simple steps:
-Step 1: Create a folder named PriceFM, along with subfolders Data, Model, and Result. Place the energy data EU_Spatiotemporal_Energy_Data.csv into Data, and place PriceFM.py inside the PriceFM folder.
+Step 1: Create a folder named PriceFM, along with subfolders Data, Figure, Model, and Result.
+Place the energy data EU_Spatiotemporal_Energy_Data.csv into Data, and place PriceFM.py inside the PriceFM folder.
 Step 2: Run Main.py to process the energy data, and to train, validate, and test the PriceFM. The script PriceFM.py contains all necessary functions and classes.
-Step 3: After execution, check the Model and Result folders for the saved model weights and evaluation metrics.
-Hardware and Computation
+Step 3: After execution, you can inspect: Figure/ for visualizations of forecasts versus true prices; Model/ for saved model weights; Result/ for evaluation metrics.
+Optional: To better understand the code structure and functionality, run Tutorial.ipynb block by block.
+Appendix B Hardware and Computation
 The PriceFM is evaluated on both an NVIDIA A100 GPU and an Intel Core i7-1265U CPU, respectively.
@@
 In contrast, the Intel i7-1265U is a power-efficient CPU commonly found in standard laptops.
-This dual evaluation setup allows us to assess the feasibility of deploying our model in environments without access to advanced GPU hardware.
 Under the training setup described in Section Model, the training time is approximately 4–5 minutes on the A100 GPU and 12–13 minutes on the i7 CPU.
-Due to the small batch size used, the GPU does not achieve a significant speedup over the CPU.
 Inference time for both setups is under 10 seconds.
-We note that neither training nor inference time is critical for our application, as bid submissions can occur at any point before the market gate closure.
-Lookup Table and Feature Availability
+We note that neither training nor inference time is critical for our application, as bid submissions can occur at any point before the market gate closure on a daily basis.
+Appendix C Lookup Table and Feature Availability
 The country-region code lookup table and the feature availability are listed in Table 1.
-Country Region Code Price Load Solar Wind (Onshore) Wind (Offshore)
+Table 1: Lookup table and feature availability across European regions.
+✓ indicates that the feature is available.
+(✓) denotes partial availability, and the feature is excluded from this study due to the high rate of missing values.
+Country Region Code Price Load Solar
+Wind
+Wind
 Austria AT
@@
 Slovakia SK (✓)
-Table 1: Lookup table and feature availability across European regions.
-✓ indicates that the feature is available.
-(✓) denotes partial availability, and the feature is excluded from this study due to the high rate of missing values.
-Baselines
+Appendix D Baselines and Hyperparameters
 We compare the PriceFM with several spatial and temporal models.
@@
 The number of parameters of PriceFM is only 3.38M.
-The spatial models, such as GraphConv, GraphAttn, GraphDiff, GraphSAGE, and GraphARMA, lack an explicit spatial decay mechanism guided by graph distance.
+We use the Adam optimizer with an initial learning rate of $4\times 10^{-3}$, which decays exponentially by a factor of 0.95 every 10 epochs.
+The batch size is 8, which introduces a slight regularization effect.
+The model is trained for 50 epochs, and the checkpoint with the lowest validation loss is saved.
+The graph models, such as GCN, GAT, GraphDiff, GraphSAGE, and GraphARMA, lack an explicit spatial decay mechanism guided by graph distance.
 GraphConv assigns uniform weights to all neighbors, ignoring spatial relevance; GraphAttn learns attention weights entirely from data without structural priors; GraphDiff uses diffusion kernels that spread information globally, but the importance of nodes is still determined through learned weights, without decay constraints; GraphSAGE aggregates features from sampled neighbors but lacks a notion of spatial proximity; and GraphARMA applies recursive smoothing, which can propagate noise from irrelevant neighbors.
@@
 The hyperparameters are optimized based on validation loss, and the search space is summarized in Table 3.
-All spatial models require an adjacency matrix as input, detailed in Section Spatial Topology and Adjacency Matrix.
+All spatial models require an adjacency matrix as input, detailed in Appendix E.
 Notably, these optimized spatial models contain more than 10M parameters.
-FEDFormer, TimesNet, and TimeXer have demonstrated strong performance in general time-series forecasting tasks.
-FEDFormer and TimeXer are Transformer-based architectures, while TimesNet is CNN-based.
-However, these models are not designed for spatial graph forecasting tasks and may require large-scale data to implicitly learn spatial dependencies.
+FEDFormer, iTransformer, PatchTST, TimesNet, and TimeXer have demonstrated strong performance in general time-series forecasting tasks.
+FEDFormer, iTransformer, PatchTST, and TimeXer are Transformer-based architectures, while TimesNet is CNN-based.
+However, these models are not designed for graph forecasting tasks and may require large-scale data to implicitly learn spatial dependencies.
 The hyperparameters are optimized based on validation loss, and the search space is summarized in Table 4.
-Notably, the number of parameters of these optimized temporal models ranges from 5M to 10M, and they are limited to single-region prediction.
+Notably, the number of parameters of these optimized temporal models ranges from 4.96M to 10.31M, and they are limited to single-region prediction.
 As a result, 38 separate models must be trained for each output region, since they are not designed to support multi-region forecasting.
@@
 Model Search Space
-GraphConv
+GCN
 hidden_size: {32, 128, 512}
@@
 epochs: 50
-GraphAttn
+GAT
 hidden_size: {32, 128, 512}
@@
 epochs: 50
-Table 4: Hyperparameter search space for temporal models.
+Table 4: Hyperparameter search space for time-series foundation models.
 Model Search Space
@@
 dropout: {0.1, 0.3, 0.5}
+batch_size: {8, 32, 128}
+learning_rate: {4e-4, 1e-3, 4e-3}
+epochs: 50
+iTransformer
+hidden_size: {32, 128, 512}
+e_layers: {2, 3, 4}
+d_ff: {512, 1024, 2048}
+n_heads: {2, 4, 8}
+dropout: {0.1, 0.3, 0.5}
+batch_size: {8, 32, 128}
+learning_rate: {4e-4, 1e-3, 4e-3}
+epochs: 50
+PatchTST
+hidden_size: {32, 128, 512}
+e_layers: {2, 3, 4}
+n_heads: {2, 4, 8}
+dropout: {0.1, 0.3, 0.5}
+patch_len: {4, 6, 12}
 batch_size: {8, 32, 128}
@@
 epochs: 50
-Spatial Topology and Adjacency Matrix
+Appendix E Adjacency Matrix
 We model the European market as a graph $G=(\mathcal{R},\mathcal{E})$, where each node $r\in\mathcal{R}$ is a bidding zone and edges indicate direct power flow via cross-border interconnections.
 This spatial topology is detailed in Table 6.
-Let $\mathcal{N}(r)$ denote the set of directly connected neighbors of $r$, excluding $r$ itself. The binary adjacency matrix $A\in\{0,1\}^{|\mathcal{R}|\times|\mathcal{R}|}$ is defined by
+Let $\mathcal{N}(r)$ denote the set of directly connected neighbors of $r$, excluding $r$ itself.
+The binary adjacency matrix $A\in\{0,1\}^{|\mathcal{R}|\times|\mathcal{R}|}$ is defined by
 $A_{r,s}=\begin{cases}1,&\text{if }s\in\mathcal{N}(r),\\ 0,&\text{otherwise,}\end{cases}\quad r,s\in\mathcal{R}.$ (23)
 For GNN layers, self-loops can be added via $\tilde{A}=A+I$.
-Scaling
+Appendix F Data Scaling
 To normalize the data while being robust to extreme values, we employ a RobustScaler fitted on the training data, using the Scikit-Learn implementation.
 The fitted scaler is then used to transform validation and testing data.
-Metrics
-We evaluate the model using Q10, Q50, Q90, AQL, AQCR, RMSE, MAE, and $R^{2}$, all averaged over regions and samples.
-Q10, Q50, Q90
-We compute quantile loss separately for each target quantile. For example, the Q10 loss is:
-$\mathrm{Q10}=\frac{1}{N\left|\mathcal{R}\right|\mathcal{T}}\sum_{i=1}^{N}\sum_{r_{\mathrm{out}}\in\mathcal{R}}\sum_{t=1}^{\mathcal{T}}L_{0.1}\left(y_{i,r_{\mathrm{out}},t},\hat{y}_{i,r_{\mathrm{out}},t,0.1}\right),$ (27)
-and similarly for Q50 and Q90 with $\tau=0.5$ and $0.9$.
-AQL
+Appendix G Metrics
+G.1 Quantile Loss at Individual Levels
+We compute quantile loss separately for each target quantile:
+$\text{Q}_{\tau}=\frac{1}{N\left|\mathcal{R}\right|\mathcal{T}}\sum_{i=1}^{N}\sum_{r_{\mathrm{out}}\in\mathcal{R}}\sum_{t=1}^{\mathcal{T}}L_{\tau}\left(y_{i,r_{\mathrm{out}},t},\hat{y}_{i,r_{\mathrm{out}},t,\tau}\right),$ (24)
+where $\tau\in\{0.1,0.5,0.9\}$.
+G.2 Average Quantile Loss (AQL)
 AQL represents the average quantile loss across all quantiles, as described in Section Model.
-AQCR
-AQCR captures the proportion of forecasted distributions that violate quantile monotonicity, i.e., when a higher quantile is predicted to be smaller than a lower one.
-The quantile crossing indicator for a quantile pair $(\tau_{l},\tau_{u})$ with $\tau_{l}<\tau_{u}$ is defined as:
-$C_{\tau_{l},\tau_{u}}(i,r,t)=\mathbb{1}\left[\hat{y}_{i,r,t,\tau_{l}}>\hat{y}_{i,r,t,\tau_{u}}\right],$ (28)
-where $\mathbb{1}[\cdot]$ is the indicator function. Then, the AQCR is given by:
-$\mathrm{AQCR}=\frac{1}{N\left|\mathcal{R}\right|\mathcal{T}}\sum_{i=1}^{N}\sum_{r\in\mathcal{R}}\sum_{t=1}^{\mathcal{T}}\mathrm{Cross}_{i,r,t}.$ (29)
-RMSE
+G.3 Average Quantile Crossing Rate (AQCR)
+AQCR captures the proportion of forecasted distributions that violate quantile monotonicity, i.e., when a lower quantile is predicted to be greater than a higher one.
+For each sample, the quantile crossing indicator is defined as:
+$C_{i,r,t}=\mathbb{I}\left(\max_{\tau_{l}<\tau_{u}}\left(\hat{y}_{i,r,t,\tau_{l}}-\hat{y}_{i,r,t,\tau_{u}}\right)>0\right)$ (25)
+where $\mathbb{I}(\cdot)$ is an indicator function that returns 1 if any quantile pair fulfills the condition inside and 0 otherwise.
+We compute the AQCR as:
+$\mathrm{AQCR}=\frac{1}{N\left|\mathcal{R}\right|\mathcal{T}}\sum_{i=1}^{N}\sum_{r\in\mathcal{R}}\sum_{t=1}^{\mathcal{T}}C_{i,r,t}.$ (26)
+A lower AQCR indicates fewer quantile crossing violations and thus reflects more reliable probabilistic forecasts.
+G.4 Root Mean Squared Error (RMSE)
 We compute RMSE within each region, then average over all regions:
@@
 $\mathrm{RMSE}=\frac{1}{\left|\mathcal{R}\right|}\sum_{r\in\mathcal{R}}\mathrm{RMSE}_{r}.$ (28)
-MAE
+G.5 Mean Absolute Error (MAE)
 Same procedure as RMSE, but using absolute error:
@@
 $\mathrm{MAE}=\frac{1}{\left|\mathcal{R}\right|}\sum_{r\in\mathcal{R}}\mathrm{MAE}_{r}.$ (30)
-$R^{2}$
-We compute $R^{2}$ for each region and average across all regions:
-$R^{2}_{r}=1-\frac{\sum_{i=1}^{N}\sum_{t=1}^{\mathcal{T}}\left(y_{i,r,t}-\hat{y}_{i,r,t,0.5}\right)^{2}}{\sum_{i=1}^{N}\sum_{t=1}^{\mathcal{T}}\left(y_{i,r,t}-\bar{y}_{r}\right)^{2}},$ (34)
+G.6 Coefficient of Determination
+We compute the Coefficient of Determination ($\text{R}^{2}$) for each region and average across all regions:
+$\text{R}^{2}_{r}=1-\frac{\sum_{i=1}^{N}\sum_{t=1}^{\mathcal{T}}\left(y_{i,r,t}-\hat{y}_{i,r,t,0.5}\right)^{2}}{\sum_{i=1}^{N}\sum_{t=1}^{\mathcal{T}}\left(y_{i,r,t}-\bar{y}_{r}\right)^{2}},$ (31)
 $\bar{y}_{r}=\frac{1}{N\mathcal{T}}\sum_{i=1}^{N}\sum_{t=1}^{\mathcal{T}}y_{i,r,t},$ (32)
-$R^{2}=\frac{1}{\left|\mathcal{R}\right|}\sum_{r\in\mathcal{R}}R^{2}_{r}.$ (36)
-DM Test
+$\text{R}^{2}=\frac{1}{\left|\mathcal{R}\right|}\sum_{r\in\mathcal{R}}\text{R}^{2}_{r}.$ (33)
+G.7 Diebold & Mariano (DM) Test
 To assess whether differences in forecasting performance are statistically significant, we apply the DM test. We compute the loss differential at each prediction instance.
@@
 The index $j$ enumerates all prediction instances across dimensions, and $\hat{\sigma}_{d}$ is the sample standard deviation of $\{d_{j}\}_{j=1}^{M}$.
-We compute a two-sided $p$-value using the Student’s $t$-distribution with M1M-1 degrees of freedom.
-A significantly positive (negative) DM value indicates that model 2 (model 1) significantly outperforms the other.
+We compute a $p$-value; if $p<0.05$ and the DM value is positive (negative), then we report that model 2 (model 1) significantly outperforms the other in Section Experiment.
+The rules are summarized in Table 5.
+Table 5: Interpretation of DM test outcomes.
+Condition Interpretation Conclusion
+$p<0.05$, DM >0>0 Statistically significant Model 2 is better
+$p<0.05$, DM <0<0 Statistically significant Model 1 is better
+$p\geq 0.05$ Not statistically significant
 Table 6: Direct neighbors by region and neighbor count.